### 12 Jan 2024 v2

# Involutive Markov categories and the quantum de Finetti theorem

Abstract
Markov categories have recently emerged as a powerful high-level framework for probability theory and theoretical statistics. Here we study a quantum version of this concept, called involutive Markov categories. First, we show that these are equivalent to Parzygnat’s quantum Markov categories but argue that they are simpler to work with. Our main examples of involutive Markov categories involve C*-algebras (of any dimension) as objects and completely positive unital maps as morphisms in the picture of interest. Second, we prove a quantum de Finetti theorem for both the minimal and the maximal C*-tensor norms, and we develop a categorical description of such quantum de Finetti theorems which amounts to a universal property of state spaces.
1 Introduction
Categorical probability, and in particular the theory of Markov categories, has attracted much interest in recent years. This ‘‘synthetic’’ approach to probability has proven fruitful in several areas, including statistics [17], graphical models [19] and ergodic theory [35]. The first paper on a quantum version of Markov categories was written by Parzygnat [38], whose main topic has been information flow axioms (see [15] on this term) and Bayesian inversion. Although there are many examples of ordinary Markov categories [13], the only quantum Markov category considered so far has finite-dimensional C*-algebras as objects. Roughly speaking, a C*-algebra is an algebra over the complex numbers, equipped with an involution modelled after hermitian conjugation and a norm which makes it into a Banach space. [^1] In contrast to most references on C*-algebras, we will always assume our C*-algebras to be unital. In quantum physics, C*-algebras come up in the form of algebras of observables. In this context, physical transformations or processes are modelled as completely positive unital maps between C*-algebras. Probabilistic Gelfand duality [22] (Proposition A.14) then explains the sense in which this formalism is a noncommutative generalization of probability: The category of commutative C*-algebras with positive unital maps is dually equivalent to the category of compact Hausdorff spaces with continuous Markov kernels.
The present paper develops a quantum version of Markov categories which enables a sensible treatment of infinite-dimensional C*-algebras. Among other things, this results in a new perspective on the quantum de Finetti theorem in terms of a universal property of state spaces. To the best of our knowledge, the version of the quantum de Finetti that we prove strengthens the known results in the theory of operator algebras.
Quantum versions of Markov categories.
A first important step is the definition of involutive Markov categories. These categories are equivalent to Parzygnat’s quantum Markov categories (Theorem 2.6), but they avoid any distinction between ‘‘even’’ and ‘‘odd’’ morphisms—which correspond to linear and antilinear maps, respectively—thus making the theory a bit cleaner. Another helpful notion introduced in this paper is that of picture, which encodes some requirements one would expect to have when dealing with quantum probability. For instance, the classical objects of any picture—i.e., those conforming to classical probability—form a Markov category. To better understand this concept, let us consider the example of C*-algebras. The associated involutive Markov category contains all unital maps, but only the completely positive ones have a well-behaved and physically relevant probabilistic interpretation. Therefore, we consider the picture given by the completely positive unital maps, and keep the involutive Markov category as the environment for diagrammatic calculus.
As pointed out by Parzygnat [38, Remark 3.12], a major challenge for quantum probability is the construction of a quantum Markov category of infinite-dimensional C*-algebras. The issue is that the multiplication map on an infinite-dimensional C*-algebra is typically unbounded [38, Remark 3.12]. Since the multiplication maps are the copy morphisms, which form an essential piece of structure in any flavour of Markov category, we are forced to allow unbounded maps as morphisms. In order for these to compose reasonably, we make their domains part of the objects under consideration by working with C*-pairs, where a C*-pair is a C*-algebra equipped with a dense ***-subalgebra (Definition 2.27). Like this, we obtain a well-behaved involutive Markov category containing infinite-dimensional C*-algebras, and with completely positive unital maps as the picture of interest.
Another issue with considering infinite-dimensional C*-algebras is that there is no canonical tensor product for them, as there are two generally distinct canonical norms for the tensor product of two C*-algebras. These are known as the minimal and maximal C*-tensor norms. In particular, the choice of either gives rise to different pictures and involutive Markov categories. Although all our results apply to both cases, the distinction between the two norms also manifests itself at the categorical level: see Proposition 3.12.
The quantum de Finetti theorem.
In classical probability, the de Finetti theorem tells us that whenever the probability of an infinite sequence of outcomes is independent of their order, such outcomes are drawn independently from the same underlying distribution (which can itself be random). Historically, this result has been a milestone in the debate over the subjective vs. objective view of probability, but it is also crucial to the development of nonparametric Bayesian modeling. A brief discussion can be found in [18], where the authors give a new proof of the de Finetti theorem using only the technology of Markov categories. Furthermore, this synthetic approach generalizes the result by allowing everything to depend measurably on an additional parameter.
A quantum analog of the de Finetti theorem is due to Størmer [43], and it is fundamental in quantum Bayesianism [21]. A categorical perspective on the quantum de Finetti theorem has already been considered by Staton and Summers [42], who extended Størmer’s result with parameters using categorical limits. There is, however, one major drawback: These works focus only on the minimal C*-tensor norm. The case of the maximal C*-tensor norm was considered by Hulanicki and Phelps [25], but the paper seems to be unknown. Their argument, which holds for both the minimal and the maximal C*-tensor norm, is presented in Appendix C. Moreover, this generalizes to the case of completely positive unital maps with an additional tensor factor (Theorem 3.69), in particular allowing for correlation and (a priori) entanglement with another system. Our result then in particular implies that there is no entanglement with the additional system, a fact whose finite-dimensional special case is well-known in quantum information theory, as it underlies a prominent separability criterion for quantum states [10, Theorem 1].
In our categorical setting, this quantum de Finetti theorem (for either choice of tensor product) results in an elegant universal property similar to the one of Staton and Summers [42]. This limit is a universal property of the state space of a C*-algebra, and it gives rise to a quantum analog of the notion of representability of a Markov category [17].
Overview
Section 2 focuses on involutive Markov categories. The definition, together with a discussion of the notion and the relation to that of quantum Markov categories, is covered in Section 2.1. We then study the relevant notions of functors and natural transformations leading to a strictification theorem for such categories (Sections 2.2 and 2.3). A first example of involutive Markov categories is given in Section 2.4, while Section 2.5 introduces the category of C*-pairs, which will be used throughout. Sections 2.6 and 2.7 develop basic properties of objects and morphisms: The former discusses new key notions, such as classicality of objects and compatibility of morphisms, while the latter defines four non-equivalent definitions of almost sure equality, all of which specialize to the standard one for classical Markov categories.
Our main results appear in Section 3. We begin by introducing pictures (Section 3.1) and extend the characterization of almost sure equality developed by Parzygnat [38, Section 5] (Section 3.2). The following subsections present quantum versions of notions that already appear in the theory of Markov categories: Kolmogorov products [20] in Section 3.3 and representability [17] in Sections 3.4 and 3.5. Section 3.6 is devoted to the quantum de Finetti theorem in its categorical formulation, which relies on the concepts developed up to that point.
In Appendix A, we present some well-known results from the theory of C*-algebras and extend them to the setting of pre-C*-algebras. Appendix B supplies a proof omitted from Remark 3.29 concerning the Hewitt–Savage zero–one law. The quantum de Finetti theorem for states is discussed in Appendix C.
Prerequisites
In this article, we will make extensive use of string diagrams for symmetric monoidal categories. An introduction to such diagrams can be found in [41]. We point out that this graphical approach suppresses the coherence isomorphisms of the monoidal structure. For this reason, it is important to understand if the categories we work with admit a strictification. This turns out to be the case for our context—a precise statement is Theorem 2.15.
Besides string diagrams, the article is intended to be readable by people with a basic background in symmetric monoidal categories, although some knowledge of existing work on Markov categories or quantum Markov categories [38, 13] may improve the understanding of the paper.
For the sake of completeness, all necessary basics about (pre-)C*-algebras are covered in Appendix A, but a previous acquaintance with these algebras may help parsing some arguments.
Acknowledgements
We thank Tomáš Gonda, Paolo Perrone, Sam Staton, Ned Summers and Nico Wittrock for fruitful discussions and comments. The authors acknowledge support from the Austrian Science Fund (FWF P 35992-N).
2 Involutive Markov categories
In this section, we introduce the protagonists of our approach to categorical quantum probability, discuss some basic examples and properties, and prove a strictification theorem.
We begin by introducing involutive Markov categories, and the more general ICD-categories, as equivalent reformulations of Parzygnat’s notions of quantum Markov category and quantum CD-category, respectively. The main difference is that our categories do not involve a distinction between ‘‘even’’ and ‘‘odd’’ morphisms, since they only contain the even ones to start with. This makes them somewhat simpler to work with than Parzygnat’s original concepts. As the names indicate, our categories have an involution on morphisms instead. We then state and prove a strictification theorem for ICD-categories (Theorem 2.15). We also introduce a notion of involutive comonoid in a hom-involutive symmetric monoidal category such that the category of all such comonoids becomes an involutive Markov category (Proposition 2.19). In particular, this answers the last question in [38, Question 3.24]. Then we introduce our main example, the involutive Markov category of C*-pairs, where a C*-pair is a C*-algebra together with a dense ***-subalgebra in it. This choice is motivated by the fact that there is no quantum Markov category of all C*-algebras [38, Remark 3.12], and therefore we need to adapt the framework to something slightly different. Basically, C*-pairs have a better behaved tensor product, and this allows us to consider the multiplication as a morphism despite it generally being unbounded. The section concludes by discussing some important notions of compatibility of morphisms related to compatibility of measurements in quantum theory, and by specializing almost sure equalities to our context.
2.1 Definition and relation to quantum Markov categories
Throughout, we work in a symmetric monoidal category $({\mathsf{C}},\otimes,I)$ together with a strict symmetric monoidal endofunctor
$\mathfrak{I}\colon{\mathsf{C}}\to{\mathsf{C}}$
called involution, such that $\mathfrak{I}^{2}=\mathrm{id}_{{\mathsf{C}}}$ and $\mathfrak{I}{(A)}=A$ for every object $A$ of ${\mathsf{C}}$. When needed, such a ${\mathsf{C}}$ will be called hom-involutive symmetric monoidal category. [^2] Alternatively, this is a symmetric monoidal category enriched in $\mathbb{Z}_{2}$-sets. We leave the details of the equivalence to the interested reader. We will write $\mathfrak{I}{(\phi)}=\overline{\phi}$ for morphisms.
Definition 2.1.
A hom-involutive symmetric monoidal category ${\mathsf{C}}$ is an ICD-category if every object $A$ comes equipped with a copy morphism $\operatorname{copy}_{A}\colon A\to A\otimes A$ and a delete morphism $\operatorname{del}_{A}:A\to I$, diagrammatically written as
[…]
such that:
  1. (i)
    Every object becomes a comonoid:
[…]
  2. (ii)
    They behave nicely with the involution:
[…]
    where we use $\overline{\bullet}$ to denote $\overline{\operatorname{copy}_{A}}$ and $\overline{\operatorname{del}_{A}}$, respectively.
  3. (iii)
    They respect the symmetric monoidal structure:
[…]
As is commonly done in string diagram calculus, we leave out the associators and unitors in the diagrams. For example, the final equation in (2.3) needs to be understood as holding with respect to the coherence isomorphism $I\otimes I\stackrel{{\scriptstyle\cong}}{{\to}}I$.
Definition 2.2.
A morphism $\phi\colon A\to B$ in an ICD-category ${\mathsf{C}}$ is
  1. (i)
    total if $\operatorname{del}_{B}\mathchoice{\,}{\,}{}{}\phi=\operatorname{del}_{A}$, i.e.
[…]
  2. (ii)
    self-adjoint if $\phi=\overline{\phi}$.
If all morphisms in ${\mathsf{C}}$ are total, ${\mathsf{C}}$ is called an involutive Markov category.
In particular, all associators, unitors and swap morphisms are self-adjoint as the involution functor $\mathfrak{I}$ is assumed strict symmetric monoidal. Similarly, identities are self-adjoint as $\mathfrak{I}$ is identity-on-objects.
Remark 2.3.
An involutive Markov category is, equivalently, an ICD-category whose monoidal unit is terminal. In fact, delete morphisms are natural precisely on total morphisms.
Remark 2.4 (ICD-categories and involutive symmetric monoidal categories).
Symmetric monoidal categories equipped with an involution have already appeared in the literature: see, for instance, [26, Definition 4]. Although our ICD-categories are involutive symmetric monoidal categories (with additional structure), our conditions on the involution are stronger than what is generally required in that context: our requirement $A=\overline{A}$ for all objects $A$ is not satisfied for typical involutive symmetric monoidal categories. A good example is the category of complex vector spaces, where $\overline{A}$ for a vector space $A$ is the complex conjugate vector space, which has the same vectors as $A$ but conjugate scalar multiplication,
$\lambda\cdot_{\overline{A}}a\coloneqq\overline{\lambda}\cdot_{A}a.$
Clearly this category does not respect the equality $A=\overline{A}$ required for hom-involutive symmetric monoidal categories. ICD-categories with C*-algebras as objects will be our main examples. It is conceivable that using complex conjugation as involution on C*-algebras could result in an interesting approach different from ours [38, Question 3.24]. This amounts to equipping every object $A$ with an involution morphism *$*_{A}\colon A\to\overline{A}$. Instead of going in this direction, we prefer to avoid fixing such a star operation [^3] We prefer to reserve the word “involution” for our functor $\mathfrak{I}\colon\phi\mapsto\overline{\phi}$ and therefore refer to the operation *** on a C*-algebra or similar structure as “star”. for every object, since this results in a more complex setting and we currently see no real advantage in such additional complexity. For our purposes, it is enough that the axioms fix how the involution acts on copy and delete morphisms, which is exactly what we have in Definition 2.1, and allows us to talk about self-adjoint morphisms. In our examples with C*-algebras as objects, the involution of a morphism will be defined in terms of conjugation by *** as we will consider it in (2.4) and (2.9). A potential drawback of this approach is that we cannot use the star operation itself in string diagrams, since it is not even a morphism in our categories of interest. This issue will be addressed by Theorem 2.6.
We will restrict our focus to involutive Markov categories from Section 2.5 onward. For the time being, let us state some important connections between ICD-categories and the existing theory of CD-categories [7, 13] and quantum CD-categories in the sense of Parzygnat [38]. We will then explain how our definition better unifies the classical and quantum cases.
Remark 2.5.
  1. (i)
    CD-categories are precisely the ICD-categories with trivial involution.
  2. (ii)
    To any quantum CD-category ${\mathsf{C}}^{\prime}$ [38, Definition 3.4], we can associate an ICD-category by restricting to even morphisms and considering the involution
    $\overline{\phi}\coloneqq*_{B}\mathchoice{\,}{\,}{}{}\phi\mathchoice{\,}{\,}{}{}*_{A}$ (2.4)
    for every morphism $\phi\colon A\to B$. The ICD-category so obtained is denoted $J({\mathsf{C}}^{\prime})$. Then $\phi$ is self-adjoint in $J({\mathsf{C}}^{\prime})$ if and only if it is $\ast$-preserving in the sense of [38, Definition 3.15] in ${\mathsf{C}}^{\prime}$.
Theorem 2.6.
The following two kinds of structures are equivalent:
  1. (i)
    Quantum CD-categories;
  2. (ii)
    ICD-categories.
More precisely, there is a construction ${\mathsf{C}}\mapsto Q({\mathsf{C}})$ turning every ICD-category ${\mathsf{C}}$ into a quantum CD-category $Q({\mathsf{C}})$ such that:
  1. (i)
    $JQ({\mathsf{C}})={\mathsf{C}}$ for every ICD-category ${\mathsf{C}}$;
  2. (ii)
    $QJ({\mathsf{C}}^{\prime})$ is canonically strictly monoidally isomorphic to ${\mathsf{C}}^{\prime}$ for any quantum CD-category ${\mathsf{C}}^{\prime}$.
Proof.
We already described the ICD-category $J({\mathsf{C}}^{\prime})$ associated to a quantum CD-category ${\mathsf{C^{\prime}}}$ in Remark 2.5(ii). Let us now explain in full detail how we get a quantum CD-category $Q({\mathsf{C}})$ from an ICD-category ${\mathsf{C}}$. This $Q({\mathsf{C}})$ is defined as follows:
  • Its objects are the ones of ${\mathsf{C}}$.
  • Its hom-sets are
    ${Q({\mathsf{C}})}(A,B)\coloneqq{{\mathsf{C}}}(A,B)\times\{0,1\}.$
    To simplify notation, we will write $\phi$ instead of $(\phi,0)$ and $\phi^{*}$ instead of $(\phi,1)$. The former are the even morphisms in $Q({\mathsf{C}})$ while the latter are the odd ones.
  • Composition is given by the following rules, where the left-hand side is composition in $Q({\mathsf{C}})$ and the right-hand side is composition in ${\mathsf{C}}$:
    $\psi\mathchoice{\,}{\,}{}{}\phi\coloneqq\psi\mathchoice{\,}{\,}{}{}\phi,\qquad\psi\mathchoice{\,}{\,}{}{}\phi^{*}\coloneqq(\psi\mathchoice{\,}{\,}{}{}\phi)^{*},\qquad\psi^{*}\mathchoice{\,}{\,}{}{}\phi\coloneqq\psi\mathchoice{\,}{\,}{}{}\overline{\phi}^{*},\qquad\psi^{*}\mathchoice{\,}{\,}{}{}\phi^{*}\coloneqq\psi\mathchoice{\,}{\,}{}{}\overline{\phi}.$
    In particular, we have $\phi^{*}=\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$, and the equations are then constructed precisely such that $\overline{\phi}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}$. Since $\mathrm{id}^{*}\mathrm{id}^{*}=\mathrm{id}$, this also shows that an even morphism $\phi$ is self-adjoint in ${\mathsf{C}}$ in our sense if and only if $\phi\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\phi$, i.e. $\phi$ is ***-preserving.
  • The graded symmetric monoidal structure is such that the tensor product of even morphisms is the one of ${\mathsf{C}}$, while we define the tensor of odd morphisms as
    $\phi^{*}\otimes\psi^{*}\coloneqq(\phi\otimes\psi)^{*}.$ (2.5)
    The associators, unitors and swap morphisms are those of ${\mathsf{C}}$ itself included in $Q({\mathsf{C}})$ as even morphisms, where the relevant coherences hold simply because they hold in ${\mathsf{C}}$. For example, let us verify naturality of the associator $\alpha$ with respect to odd morphisms. By the definition above,
    $\alpha\mathchoice{\,}{\,}{}{}(\phi^{*}\otimes(\psi^{*}\otimes\omega^{*}))=\alpha\mathchoice{\,}{\,}{}{}(\phi^{*}\otimes(\psi\otimes\omega)^{*})=\alpha\mathchoice{\,}{\,}{}{}(\phi\otimes(\psi\otimes\omega))^{*}=((\phi\otimes\psi)\otimes\omega)\mathchoice{\,}{\,}{}{}\alpha\mathchoice{\,}{\,}{}{}\mathrm{id}^{*},$
    so that the problem is reduced to showing $\alpha\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}=\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\alpha$, which holds because $\alpha$ is self-adjoint in ${\mathsf{C}}$. Similar arguments show naturality of the unitors and the swap morphisms.
We now show that $Q({\mathsf{C}})$ is a quantum Markov category with respect to *$**_{A}\coloneqq\mathrm{id}_{A}^{*}$ on every object $A$. First of all, we note $*\overline{\phi}=*\mathchoice{\,}{\,}{}{}\phi\mathchoice{\,}{\,}{}{}*$ for all morphisms $\phi$. We then prove that Parzygnat’s (QCD3) axiom holds:
  • $\mathrm{id}^{*}\mathchoice{\,}{\,}{}{}\mathrm{id}^{*}=\mathrm{id}\mathchoice{\,}{\,}{}{}\overline{\mathrm{id}}=\overline{\mathrm{id}}=\mathrm{id}$,
  • $\mathrm{id}_{A\otimes B}^{*}=(\mathrm{id}_{A}\otimes\mathrm{id}_{B})^{*}=\mathrm{id}_{A}^{*}\otimes\mathrm{id}_{B}^{*}$.
  • Moreover, $\operatorname{del}\mathchoice{\,}{\,}{}{}*=*\mathchoice{\,}{\,}{}{}\operatorname{del}$ because $\operatorname{del}=\overline{\operatorname{del}}$.
The only thing we still need to prove is the last condition of Parzygnat’s (QCD1), i.e.
$\operatorname{copy}\mathchoice{\,}{\,}{}{}*=\operatorname{swap}\mathchoice{\,}{\,}{}{}*\mathchoice{\,}{\,}{}{}\operatorname{copy}\!.$
This follows from
$\operatorname{copy}\mathchoice{\,}{\,}{}{}*=*\mathchoice{\,}{\,}{}{}*\mathchoice{\,}{\,}{}{}\operatorname{copy}\mathchoice{\,}{\,}{}{}*=*\mathchoice{\,}{\,}{}{}\overline{\operatorname{copy}}=*\mathchoice{\,}{\,}{}{}\operatorname{swap}\mathchoice{\,}{\,}{}{}\operatorname{copy}=\operatorname{swap}\mathchoice{\,}{\,}{}{}*\mathchoice{\,}{\,}{}{}\operatorname{copy},$
where in the last equality we used the self-adjointness of the swap morphism. From this construction, we immediately conclude that $JQ({\mathsf{C}})={\mathsf{C}}$ holds by definition. Conversely, starting from a quantum CD-category ${\mathsf{C}}^{\prime}$, note that all odd morphisms $\psi$ are of the form $*\phi\mathchoice{\,}{\,}{}{}*$ where $\phi$ is an even morphism. Indeed, it suffices to take $*\phi\coloneqq\psi\mathchoice{\,}{\,}{}{}*$ since **$*\mathchoice{\,}{\,}{}{}*=\mathrm{id}$. More precisely, we have a bijection
${{\mathsf{C}}^{\prime}}(A,B)^{\text{even}}\cong{{\mathsf{C}}^{\prime}}(A,B)^{\text{odd}}$
given by precomposition with ***. This implies that the even morphisms completely describe the morphisms of ${\mathsf{C}}^{\prime}$, in the sense that there is a canonical isomorphism $F$ of categories between ${\mathsf{C}}^{\prime}$ and $QJ({\mathsf{C}}^{\prime})$ given by the identity on objects and on even morphisms, and on odd morphisms the composite bijection
${{\mathsf{C}}^{\prime}}(A,B)^{\text{odd}}\cong{J({\mathsf{C}}^{\prime})}(A,B)\times\{1\}={QJ({\mathsf{C}}^{\prime})}(A,B)^{\text{odd}}$
for any two objects $A$ and $B$. Furthermore, the graded symmetric monoidal structure on $QJ({\mathsf{C}}^{\prime})$ is exactly the one induced by transporting the one on ${\mathsf{C}}^{\prime}$ via $F$. Indeed, let $\phi^{*}$ and $\psi^{*}$ be any odd morphisms in ${\mathsf{C}}^{\prime}$, then
[…]
where ()($\clubsuit$)( ♣ ) holds by (QCD3). Therefore, (2.5) still holds. As everything else is already fixed, from a direct check we conclude that $F$ is a strict symmetric monoidal isomorphism ${\mathsf{C}}^{\prime}\to QJ({\mathsf{C}}^{\prime})$, where the preservation of the coherences is obvious as these are even. ∎
Remark 2.7.
  1. (i)
    Theorem 2.6 tells us that quantum CD-categories and ICD-categories are different ways to talk about the same concept. This raises the question of which of these two formulations is preferable, if any. We think that ICD-categories should be preferred for defining and working with concrete examples, since one only needs to consider even morphisms. On the other hand, quantum CD-categories facilitate more general diagrammatic proofs by allowing the star *** to appear in string diagrams, as long as one is careful not to tensor an even with an odd morphism. For example, for every morphism $\phi\colon A\to B$ in an ICD-category, we can write
[…]
    where the circled *** denotes the star of the associated quantum CD-category. The problem with this approach is that one needs to be careful not to tensor an even with an odd morphism, as this is not defined. Moreover, for the purposes of this paper, using the star *** in string diagrams is not needed.
  2. (ii)
    While every CD-category also gives rise to a quantum CD-category, it seems unlikely that its odd morphisms will be of any interest at all. This suggests that ICD-categories encode classical examples better than quantum CD-categories do, and therefore should be slightly preferred. For example, consider the Markov category $\mathsf{FinStoch}$ with finite sets as objects and stochastic matrices as morphisms [13, Example 2.5]. Then $Q(\mathsf{FinStoch})$ can also be described as the quantum Markov category whose odd morphisms are matrices with nonpositive entries such that any column sums to $-1$ with *$*_{A}\coloneqq-\mathrm{id}_{A}$.
2.2 Deterministic morphisms and ICD-functors
Now that ICD-categories are defined, it is natural to ask what the corresponding notion of functor between such categories is. For our framework, the introduction of such a concept is actually crucial, because string diagrams do not illustrate associators and unitors: We are therefore allowed to use them as usual only if we have a strictification theorem.
In this functorial study, an important role is played by deterministic morphisms. As we will see in this section, coherence morphisms of ICD-functors turn out to be deterministic (Remark 2.12), while ICD-natural transformations are defined as having deterministic components. For this reason, we first define what a deterministic morphism is.
Definition 2.8.
Let ${\mathsf{C}}$ be an ICD-category. A deterministic morphism is a self-adjoint total morphism $\phi\colon A\to B$ such that
[…]
The subcategory of all deterministic morphisms will be denoted ${\mathsf{C}}_{\mathrm{det}}$.
Example 2.9.
Associators, unitors, and swap morphisms are deterministic, the proof being analogous to the one for classical Markov categories [13, Lemma 10.12]. The same is true for $\operatorname{del}$ simply because $\operatorname{copy}$ is total. In particular, ${\mathsf{C}}_{\mathrm{det}}$ is a semicartesian [^4] Recall that this means that its monoidal unit is terminal. symmetric monoidal subcategory of ${\mathsf{C}}$, since tensor products and composites of deterministic morphisms are easily seen to be deterministic. This generalizes the known statement for Markov categories [13, Remark 10.13] and was already briefly pointed out by Parzygnat [38, Remark 3.19] for quantum Markov categories, although the question of determinism of the coherence isomorphisms was not addressed.
An important difference between the classical and the quantum setting is that a copy morphism $\operatorname{copy}_{A}$ is deterministic if and only if $A$ is a classical object in the sense of Definition 2.36, which we will discuss in Section 2.6.
We turn our attention to functors. As for monoidal categories in general, it may be of interest to consider different versions like strong, lax or oplax functors between ICD-categories. But for our purposes, the following strong version will be sufficient.
Definition 2.10.
Let ${\mathsf{C}}$ and ${\mathsf{D}}$ be hom-involutive symmetric monoidal categories. A hom-involutive strong symmetric monoidal functor is a strong symmetric monoidal functor $F\colon{\mathsf{C}}\to{\mathsf{D}}$ such that
  1. (i)
    $F$ strictly commutes with the involution: $F(\overline{\phi})=\overline{F(\phi)}$ for every morphism $\phi$ in ${\mathsf{C}}$;
  2. (ii)
    The coherence isomorphisms $\phi_{A,B}\colon FA\otimes FB\to F(A\otimes B)$ and $\phi_{I}\colon I\to F(I)$ are self-adjoint.
This arises from the general notion of involutive monoidal functor between involutive monoidal categories [26, Definition 5] if one specializes to our setting of identity-on-objects involutions, in which case it is natural to consider the coherence isomorphisms $F(\overline{A})\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\overline{F(A)}$ to be identities as well.
Definition 2.11.
Let ${\mathsf{C}}$ and ${\mathsf{D}}$ be ICD-categories. Then $F\colon{\mathsf{C}}\to{\mathsf{D}}$ is a strong ICD-functor if
  1. (i)
    $F$ is a hom-involutive strong symmetric monoidal functor;
  2. (ii)
    The coherence isomorphisms $\phi_{A,B}\colon FA\otimes FB\to F(A\otimes B)$ and $\phi_{I}\colon I\to F(I)$ additionally satisfy
[…]
If ${\mathsf{C}}$ and ${\mathsf{D}}$ are CD-categories, then this notion specializes to the existing notion of strong CD-functor (called strong gs-monoidal functor in [14, Definition 2.5]).
Remark 2.12.
The coherence morphisms of a strong ICD-functor are necessarily deterministic. This statement is an instance of [12, Theorem 4.7].
For the relevant notion of natural transformation, see also [13, Definition 10.14] in the Markov categories context.
Definition 2.13.
Given two strong ICD-functors $F,G\colon{\mathsf{C}}\to{\mathsf{D}}$, an ICD-natural transformation $\eta\colon F\to G$ is a monoidal natural transformation whose components are deterministic.
It is then straightforward to see that ICD-categories, strong ICD-functors and ICD-natural transformations form a 2-category. The invertible 2-cells in this 2-category, which we call ICD-natural isomorphisms, are precisely those ICD-natural transformations whose components are deterministic isomorphisms. This explains why we require the components of ICD-natural transformations to be deterministic: An ICD-natural isomorphism should make the comonoid structures on any two objects between which it interpolates match up.
The equivalences in the 2-category of ICD-categories are the ICD-equivalences. Therefore an ICD-equivalence is a strong ICD-functor $F\colon{\mathsf{C}}\to{\mathsf{D}}$ for which there exist a strong ICD-functor $G\colon{\mathsf{D}}\to{\mathsf{C}}$ and ICD-natural isomorphisms $\eta\colon F\mathchoice{\,}{\,}{}{}G\to\mathrm{id}_{{\mathsf{D}}}$ and $\mu\colon G\mathchoice{\,}{\,}{}{}F\to\mathrm{id}_{{\mathsf{C}}}$. As for ordinary categories and for Markov categories [13, Proposition 10.16], ICD-equivalences can also be characterized more concretely.
Proposition 2.14.
A strong ICD-functor $F\colon{\mathsf{C}}\to{\mathsf{D}}$ is an ICD-equivalence if and only if it is fully faithful and for every $B\in{\mathsf{D}}$ there exist $A\in{\mathsf{C}}$ and a deterministic isomorphism $B\to FA$.
Sketch of proof.
Whenever $F$ is an ICD-equivalence, it suffices to take an essential inverse $G$ as above and set $A\coloneqq GB$ to ensure that $F$ is an equivalence satisfying the claimed condition with $\eta^{-1}_{B}\colon B\to FGB$ a deterministic isomorphism.
Conversely, the listed conditions on $F$ immediately imply that $F$ is a symmetric monoidal equivalence. By the assumption of admitting deterministic isomorphisms $B\to FA$ and the same arguments as in [13, Proposition 10.16], we can construct a strong symmetric monoidal inverse $G$ with a natural monoidal isomorphism $\epsilon\colon F\mathchoice{\,}{\,}{}{}G\to\mathrm{id}_{{\mathsf{D}}}$ which has deterministic components. By construction, its coherence morphisms are deterministic because Remark 2.12 ensures that those of $F$ are deterministic. Moreover, $G$ also preserves the involution: since $\epsilon:F\mathchoice{\,}{\,}{}{}G\to\mathrm{id}_{{\mathsf{D}}}$ has self-adjoint components, we conclude that also $F\mathchoice{\,}{\,}{}{}G$ preserves the involution, and hence
$F(\overline{G(\phi)})=\overline{F\mathchoice{\,}{\,}{}{}G(\phi)}=F\mathchoice{\,}{\,}{}{}G(\overline{\phi})$
for every morphism $\phi$ in ${\mathsf{D}}$. From faithfulness of $F$ we thus conclude $\overline{G(\phi)}=G(\overline{\phi})$. It remains to show that $\eta\colon\mathrm{id}\to GF$ has deterministic components and that $G$ satisfies the commutative diagrams (2.7). These properties can be checked as in the proof of [13, Proposition 10.16]. ∎
2.3 Strictification
We consider an ICD-category to be strict if and only if its underlying monoidal category is strict.
Theorem 2.15 (Strictification for ICD-categories).
Every ICD-category is ICD-equivalent to a strict one.
The main difficulty in finding a correct proof is the fact that an isomorphism in an ICD-category is not necessarily deterministic [13, Remark 10.10]. We follow the same proof strategy as in the Markov category case [13, Theorem 10.17].
Sketch of proof.
Let ${\mathsf{C}}$ be an ICD-category. We first consider a strictification ${\mathsf{C}}^{\prime}_{\mathrm{det}}$ of ${\mathsf{C}}_{\mathrm{det}}$ (recall Example 2.9), and write $F_{\mathrm{det}}:{\mathsf{C}}^{\prime}_{\mathrm{det}}\to{\mathsf{C}}_{\mathrm{det}}$ for the strong symmetric monoidal equivalence. We then define an ICD-category ${\mathsf{C}}^{\prime}$ to be the category whose objects are those of ${\mathsf{C}}^{\prime}_{\mathrm{det}}$ and whose morphisms are given by
${\mathsf{C}}^{\prime}(A,B)\coloneqq{\mathsf{C}}(F_{\mathrm{det}}(A),F_{\mathrm{det}}(B)).$
This category has a natural symmetric strict monoidal structure inherited from ${\mathsf{C}}^{\prime}_{\mathrm{det}}$ (cf. [13, Theorem 10.17]). Also, the above definition of the hom-sets shows that ${\mathsf{C}}^{\prime}$ directly inherits an involution from the one of ${\mathsf{C}}$. This involution is immediately strictly monoidal, and it is a symmetric monoidal functor because the coherence morphisms of $F_{\mathrm{det}}$ are deterministic and in particular self-adjoint. Therefore, ${\mathsf{C}}^{\prime}$ becomes a hom-involutive symmetric monoidal category.
Let us name $F\colon{\mathsf{C}}^{\prime}\to{\mathsf{C}}$ the obvious functor which extends $F_{\mathrm{det}}$. This is a strong monoidal equivalence. In particular, for every object $A$ in ${\mathsf{C}}^{\prime}$, the involutive comonoid $(FA,\operatorname{copy}_{FA},\operatorname{del}_{FA})$ can be pulled to an involutive comonoid $(A,\operatorname{copy}_{A},\operatorname{del}_{A})$. Since the coherence morphisms of $F$ are deterministic with respect to ${\mathsf{C}}$ because they coincide with those of $F_{\mathrm{det}}$, this gives rise to a natural ICD-structure on ${\mathsf{C}}^{\prime}$, and furthermore $F$ becomes a strong ICD-functor. We conclude by Proposition 2.14. ∎
2.4 Cofree ICD-categories
Cofree ICD-categories are ICD-categories that enjoy a universal property (Proposition 2.21) dual to that of free groups, hence the name. As we will see shortly, the objects of this category are involutive comonoids.
Recall that a comonoid in a symmetric monoidal category ${\mathsf{C}}$ is a triple $(A,\mu_{A},\epsilon_{A})$ where $\mu_{A}\colon A\to A\otimes A$ and $\epsilon_{A}\colon A\to I$ are like $\operatorname{copy}$ and $\operatorname{del}$ in the sense that they satisfy (2.1), i.e. the comultiplication is associative and we have left and right unit laws. In the involutive setting, it is natural to impose an additional compatibility condition with the involution.
Definition 2.16.
A comonoid $(A,\mu_{A},\epsilon_{A})$ in a hom-involutive symmetric monoidal category ${\mathsf{C}}$ is involutive if (2.2) holds, i.e.
$\overline{\mu_{A}}=\operatorname{swap}_{A,A}\mathchoice{\,}{\,}{}{}\mu_{A}\qquad\text{and}\qquad\overline{\epsilon_{A}}=\epsilon_{A}.$
Example 2.17.
If ${\mathsf{C}}$ carries the trivial involution, then an involutive comonoid is just a commutative comonoid.
Definition 2.18.
For a hom-involutive symmetric monoidal category ${\mathsf{C}}$, its cofree ICD-category $\operatorname{Cof}({\mathsf{C}})$ is the category whose objects are involutive comonoids $(A,\mu_{A},\epsilon_{A})$ and morphisms are the ones in ${\mathsf{C}}$. This category becomes symmetric monoidal by lifting the tensor of ${\mathsf{C}}$ to involutive comonoids as $\epsilon_{A\otimes B}\coloneqq\epsilon_{A}\otimes\epsilon_{B}$ and
[…]
In particular, the monoidal unit is $(I,\mathrm{id}_{I},\mathrm{id}_{I})$.
In light of Theorem 2.6, the following two propositions answer Parzygnat’s [38, Question 3.24], where it was asked whether involutive comonoids form a quantum Markov category and whether the objects in a quantum Markov category are involutive comonoids in a suitable sense.
Proposition 2.19.
For any hom-involutive symmetric monoidal category ${\mathsf{C}}$, the cofree ICD-category $\operatorname{Cof}({\mathsf{C}})$ is an ICD-category, where:
  1. (i)
    The involution $\phi\mapsto\overline{\phi}$ is the one of ${\mathsf{C}}$.
  2. (ii)
    The copy and delete morphisms on each object are given by the comonoid structure.
Proof.
This follows straightforwardly from the definition of involutive comonoid and the definition of their monoidal product. ∎
We now want to show that the cofree ICD-categories contain all ICD-categories in a similar way as to how any group is contained in the symmetric group over itself. For this, let us briefly mention that an ICD-subcategory ${\mathsf{D}}$ of an ICD-category ${\mathsf{C}}$ is simply any symmetric monoidal subcategory ${\mathsf{D}}\subseteq{\mathsf{C}}$ which is closed under the involution and contains all copy and delete morphisms. It is clear that an ICD-subcategory is then an ICD-category in its own right. This is analogous to the notion of Markov subcategory [16, Definition A.1.2].
Proposition 2.20.
Let ${\mathsf{C}}$ be an ICD-category. Then ${\mathsf{C}}$ is a full ICD-subcategory of $\operatorname{Cof}({\mathsf{C}})$.
Proof.
We just consider ${\mathsf{C}}$ as a subcategory via $A\mapsto(A,\operatorname{copy}_{A},\operatorname{del}_{A})$. ∎
The cofree ICD-category enjoys also a universal property, as the name suggests. The relevant universal arrow is the forgetful functor $\operatorname{Forg}:\operatorname{Cof}({\mathsf{C}})\to{\mathsf{C}}$.
Proposition 2.21 (Universal property of the cofree ICD-category).
Let ${\mathsf{C}}$ be an ICD-category and let ${\mathsf{D}}$ be a hom-involutive symmetric monoidal category. Then, any hom-involutive strong symmetric monoidal functor $F:{\mathsf{C}}\to{\mathsf{D}}$ factors uniquely through $\operatorname{Cof}({\mathsf{D}})$: there is a commutative diagram
[…]
for a unique strong ICD-functor $\tilde{F}$.
Proof.
It is a standard fact that strong (and more generally oplax) monoidal functors send comonoids to comonoids. Applied to the canonical comonoid structure $(A,\operatorname{copy}_{A},\operatorname{del}_{A})$ on every object $A$ in ${\mathsf{C}}$, this works by defining $\tilde{F}$ as the functor
$\tilde{F}(A)\coloneqq(FA,\psi_{A,A}\mathchoice{\,}{\,}{}{}F(\operatorname{copy}_{A}),\psi_{I}\mathchoice{\,}{\,}{}{}F(\operatorname{del}_{A})),\qquad\tilde{F}(\phi)\coloneqq F(\phi),$
where $\psi_{A,A}:F(A\otimes A)\to FA\otimes FA$ and $\psi_{I}:F(I)\to I$ are the coherence morphism of $F$, with $\psi_{A,A}=\phi_{A,A}^{-1}$ and $\psi_{I}=\phi_{I}^{-1}$ relative to (2.7). By the same diagrams (2.7), it is clear that this is the only choice that makes (2.8) commute and has a chance of making $\tilde{F}$ into an ICD-functor. Hence we already obtain the uniqueness part of the claim.
For the existence, we show that $\tilde{F}$ actually lands in $\operatorname{Cof}({\mathsf{D}})$, i.e. that the comonoids $\tilde{F}(A)$ are involutive. This follows from the commutativity of the diagram
[…]
where the commutativity of the triangle is involutivity of the comonoid together with the fact that $F$ strictly commutes with the involution, while the commutativity of the square is the symmetry preservation of $F$.
Finally, $\tilde{F}$ is clearly a strong symmetric monoidal functor with respect to the same self-adjoint coherence morphisms as $F$, and it commutes with the involution because $F$ does. ∎
We now investigate (opposite) categories of ***-algebras as instances of cofree ICD-categories. First, recall that a (unital) ***-ring $R$ is a ring $R$ together with an operation *** that is:
  • Additive, $x^{*}+y^{*}=(x+y)^{*}$;
  • Unit-preserving, $11^{*}=$ 1;
  • Multiplication-reversing, $(xy)^{*}=y^{*}x^{*}$;
  • Involutive, $x^{**}=x$.
Definition 2.22 (Involutive modules).
Let $R$ be a commutative ***-ring.
  1. (i)
    An $R$-module $M$ is called involutive if it is equipped with an additive map *$*:M\to M$ such that $m^{**}=m$ and $(rm)^{*}=r^{*}m^{*}$ for all $m\in M$ and $r\in R$.
  2. (ii)
    The category of involutive modules $R\operatorname{-iMod}$ is the category whose objects are involutive modules and whose morphisms are $R$-linear maps.
In the case $R=\mathbb{C}$ with complex conjugation, *$*:M\to M$ is thus required to be antilinear, and involutive modules are then also known as ***-vector spaces.
Remark 2.23.
The category $R\operatorname{-iMod}$ comes equipped with a natural involution that fixes the objects and is given on morphisms by
$\overline{{\phi}}(x)\coloneqq{\phi}(x^{*})^{*}.$
Further, $R\operatorname{-iMod}$ is a hom-involutive symmetric monoidal category when we consider the symmetric monoidal structure given by the tensor product of $R$-modules with involution the unique antilinear extension of
$(m\otimes n)^{*}\coloneqq m^{*}\otimes n^{*}.$
Remark 2.24.
Unfolding the definitions shows that an involutive monoid [^5] Involutive monoid is just the dual notion of involutive comonoid. in $R\operatorname{-iMod}$ is the same thing as a ***-algebra over $R$. Indeed, a monoid simply gives an associative unital algebra $A$ over $R$, while the requirement of being involutive shows that $A$ is also a ***-algebra. Since the opposite category of ***-algebras over $R$ with $R$-linear maps is therefore just $\operatorname{Cof}(R\operatorname{-iMod}^{\mathrm{op}})$, we conclude that it is an ICD-category by Proposition 2.19.
In this category, it is easy to understand the notions of morphisms we already encountered.
  1. (i)
    The total morphisms are the formal opposites of unital $R$-linear maps.
  2. (ii)
    The self-adjoint morphisms are the formal opposites of $R$-linear maps which commute with the star operation, $\phi(x)^{*}=\phi(x^{*})$.
  3. (iii)
    The deterministic morphisms are the formal opposites of unital algebra homomorphisms commuting with the star operation. These are also known as ***-homomorphisms.
Example 2.25.
If $R$ is just $\mathbb{Z}$ with the trivial involution, then the involutive monoids in $\mathbb{Z}\operatorname{-iMod}$ are precisely the ***-rings. If $R$ is the complex numbers $\mathbb{C}$ with conjugation, then the involutive monoids in (,¯$(\mathbb{C},\overline{\,\cdot\,})\operatorname{-iMod}$ are exactly the complex ***-algebras in the usual sense. It follows that complex ***-algebras with linear maps form an ICD-category with respect to the formal opposites of the multiplication and the unit as the copy and delete morphisms, respectively.
Notation 2.26.
Throughout the rest of the paper, we restrict our attention to involutive Markov categories. This makes sense, since in the main categories of interest to us, the morphisms are formal opposites of unital maps, such as quantum channels. In classical Markov categories, this restriction corresponds to the normalization of probability. Note that the total morphisms in any ICD-category form an involutive Markov category. For example, the category of ***-algebras over any ***-ring with formal opposites of unital $R$-linear maps is an involutive Markov category.
This restriction to involutive Markov categories is also motivated by the recent successful research in categorical probability, including work on quantum Markov categories [38], where the restriction to total morphisms facilitates arguments that would be impossible or significantly more cumbersome for (quantum) CD-categories in general.
2.5 C*-pairs
We refer to Appendix A for some background on C*-algebras and pre-C*-algebras.
To get closer from the involutive Markov category of ***-algebras as in Remark 2.24 to quantum probability, it seems necessary to restrict to C*-algebras as objects and bounded linear maps as morphisms (or further to some subcategory thereof, like von Neumann algebras and normal completely bounded maps). However, as already noted in [38, Remark 3.12], this is not possible since the C*-algebra multiplication, which is the formal dual of our copy, is typically not bounded for infinite-dimensional C*-algebras. If we drop the boundedness requirement, then we do not have any tensor product of morphisms, since the tensor product of C*-algebras is a completion of their algebraic tensor product, and without boundedness we no longer have a natural way to extend a linear map to the completion. In order to circumvent these issues, we expand our interest to dense ***-subalgebras.
Definition 2.27.
We consider the category $\operatorname{C*-}{\mathsf{pr}}$, defined as follows.
  • Objects are C*-pairs, i.e. pairs $(A,{}_{p}{A})$, where $A$ is a unital C*-algebra and ${}_{p}{A}\subseteq A$ is a dense ***-subalgebra. For the sake of simplicity, we often denote the C*-pair just by $A$ and write ${}_{p}{A}$ to refer to the dense ***-subalgebra.
  • Morphisms $\phi\colon A\to B$ are given by formal opposites of unital linear maps, denoted as
    $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A},$
    with the obvious composition.
Remark 2.28.
  1. (i)
    The subscript $p$ in ${}_{p}{A}$ reminds us of ${}_{p}{A}$ being a pre-C*-algebra. In fact, $\operatorname{C*-}{\mathsf{pr}}$ is equivalent to the opposite category of pre-C*-algebras and linear maps (see Proposition A.3). Our description of this category in terms of pairs is motivated by two main ideas. First, it emphasizes the role of C*-algebras and makes the rest of the paper easier to follow for readers with a C*-algebra background. Second, similar definitions are commonly used in noncommutative geometry. To see why, recall that often defines geometrical objects as topological spaces equipped with extra structure (like a smooth structure). By the philosophy of Gelfand duality, this generalizes to defining a noncommutative geometrical object as a C*-algebra equipped with a suitable dense ***-subalgebra, such as a ***-subalgebra playing the role of smooth functions [11]. For example for a compact smooth manifold $M$, the algebra of smooth functions $C^{\infty}(M)$ is indeed dense in the C*-algebra of continuous functions $C(M)$ by the Stone–Weierstrass Theorem.
  2. (ii)
    The notation $\phi^{\mathrm{op}}$ has two possible interpretations. First, it indicates a switch in the direction using the terminology of opposite categories. Second, it reminds the reader that these maps follow the operator algebra direction. In physics terminology, this means that we use the Heisenberg picture.
Example 2.29.
Let $B=\mathbb{C}^{\{0,1\}}$ be the two-dimensional commutative C*-algebra whose elements are pairs of complex numbers $(b_{0},b_{1})$, multiplication is defined componentwise, and the unit is $(1,1)$. Its only dense ***-subalgebra is ${}_{p}{B}=B$ itself. For any C*-pair $(A,{}_{p}{A})$, the morphisms $\phi:A\to B$ correspond to pairs of elements $(a_{0},a_{1})\in{}_{p}{A}\times{}_{p}{A}$ with $a_{0}+a_{1}=1$—such elements are the images of $(1,0)$ and $(0,1)$ respectively. Since such a pair is uniquely determined by one component, it follows that the morphisms $A\to\mathbb{C}^{\{0,1\}}$ can be identified with the elements of ${}_{p}{A}$. The self-adjoint morphisms correspond to the self-adjoint elements, while the formal opposites of positive unital maps correspond to the elements of the unit interval $[0,1]$ in ${}_{p}{A}$ [^6] As is standard for C*-algebras, the unit interval $[0,1]\subseteq{}_{p}{A}$ is the set of elements $x\in{}_{p}{A}$ such that $0\leq x\leq 1$.
While the category of ***-algebras is most naturally equipped with the algebraic tensor product as in Remark 2.23, $\operatorname{C*-}{\mathsf{pr}}$ admits at least two symmetric monoidal structures, namely the minimal and the maximal tensor product of C*-algebras, which are associated to the minimal and maximal tensor norms, respectively.
Notation 2.30.
For the sake of clarity, from now on we will use $\odot$ to indicate the algebraic tensor product and $\otimes$ for the tensor product of C*-algebras and C*-pairs.
Definition 2.31.
We write $\operatorname{C*-}{\mathsf{pr}}_{\min}$ for the symmetric monoidal category of C*-pairs with monoidal structure given by
$(A,{}_{p}{A})\otimes(B,{}_{p}{B})\coloneqq(A\otimes_{\min}B,{}_{p}{A}\odot{}_{p}{B}),$
and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ for the symmetric monoidal category of C*-pairs with
$(A,{}_{p}{A})\otimes(B,{}_{p}{B})\coloneqq(A\otimes_{\max}B,{}_{p}{A}\odot{}_{p}{B}),$
and the induced action on morphisms in both cases.
The fact that ${}_{p}{A}\odot{}_{p}{B}$ is indeed dense in both these C*-algebras is a standard C*-algebra argument which we offer in Proposition A.5 for the sake of completeness. The trick of working with C*-pairs rather than C*-algebras is now what allows us to obtain an involutive Markov category after all.
Remark 2.32.
The categories $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ are isomorphic, since the norms do not currently play any role. However, this will change when we consider pictures (Definition 3.2). Indeed, the subcategories of interest will require morphisms to be bounded, and this is sensitive to the choice of the tensor C*-norm.
Proposition 2.33.
$\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ are involutive Markov categories with respect to the involution defined by
$\overline{\phi^{\mathrm{op}}}(x)\coloneqq\phi^{\mathrm{op}}(x^{*})^{*}$ (2.9)
as in Remark 2.23, and copy morphisms
$\operatorname{copy}_{A}^{\mathrm{op}}\>:\>{}_{p}{A}\odot{}_{p}{A}\rightsquigarrow{}_{p}{A}$
given by multiplication.
Proof.
All relevant properties follow from the fact that the category of complex ***-algebras and unital linear maps is an involutive Markov category (Example 2.25). ∎
Remark 2.34.
In both $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$, a morphism $\phi\colon A\to B$ is self-adjoint if and only if $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ satisfies $\phi^{\mathrm{op}}(x^{*})=\phi^{\mathrm{op}}(x)^{*}$ for all $x\in{}_{p}{B}$. Similarly, $\phi$ is deterministic if and only if $\phi^{\mathrm{op}}$ is a ***-homomorphism. This is in perfect analogy with Remark 2.24.
2.6 Classicality and compatibility
In our setting, an interesting question is which objects in an involutive Markov category can be considered ‘‘classical’’ or ‘‘commutative’’ and which ones display genuinely ‘‘quantum’’ or ‘‘noncommutative’’ behavior. With this in mind, we introduce classical objects, which, as the name suggests, are those that behave as in classical probability. Whenever an object $A$ is not classical, it is still relevant to understand when a pair of morphisms out of $A$ displays ‘‘classical’’ behavior. This gives rise to the notion of compatibility.
Let us start with classical objects. In ordinary Markov categories, the copy morphisms are required to be invariant under swapping the outputs. This motivates the following.
Lemma 2.35.
The following are equivalent for an object $A$:
  1. (i)
    $\overline{\operatorname{copy}_{A}}=\operatorname{copy}_{A}$.
  2. (ii)
    $\operatorname{swap}_{A,A}\mathchoice{\,}{\,}{}{}\operatorname{copy}_{A}=\operatorname{copy}_{A}$.
  3. (iii)
    $\operatorname{copy}_{A}$ is deterministic.
Proof.
The equivalence of the first two conditions is clear by the requirement $\overline{\operatorname{copy}_{A}}=\operatorname{swap}_{A,A}\mathchoice{\,}{\,}{}{}\operatorname{copy}_{A}$. The equivalence of the second and third conditions follows by a straightforward calculation as in [13, Remark 10.2] for ordinary Markov categories. ∎
Definition 2.36.
Let ${\mathsf{C}}$ be an involutive Markov category. An object $A$ in ${\mathsf{C}}$ is classical if it satisfies the equivalent conditions of Lemma 2.35.
Definition 2.37.
The classical subcategory $C\ell({\mathsf{C}})$ of an involutive Markov category ${\mathsf{C}}$ is the subcategory whose objects are classical and morphisms are self-adjoint.
It is easy to see that $C\ell({\mathsf{C}})$ is a Markov category, and in particular that if $A$ and $B$ are classical, then so is $A\otimes B$.
Example 2.38.
In the involutive Markov category given by the total morphisms of a cofree ICD-category $\operatorname{Cof}({\mathsf{C}})$, the classical subcategory consists of the commutative comonoids with self-adjoint morphisms. Similarly in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$, the classical subcategory consists of the commutative C*-pairs with self-adjoint morphisms. These statements are immediate from Lemma 2.35 since the copy morphisms are given by multiplication.
Notation 2.39.
For two morphisms $\phi,\psi$ with the same domain $A$, and $n$ any positive integer, we will use the shorthand notations
[…]
for easier writing, where the right-hand equation defines $\phi^{(n)}$ recursively starting with $\phi^{(0)}\coloneqq\operatorname{del}_{A}$.
In $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$, the left-hand equation amounts to the formation of the product map, which is defined as the unique linear extension of
$(\phi,\psi)^{\mathrm{op}}(x\odot y)\coloneqq\phi^{\mathrm{op}}(x)\,\psi^{\mathrm{op}}(y).$
Definition 2.40.
Let $\phi,\psi$ be morphisms with the same domain. Then $\phi$ and $\psi$ are compatible if $(\phi,\psi)=\operatorname{swap}\mathchoice{\,}{\,}{}{}(\psi,\phi)$. We say that $\phi$ is autocompatible if it is compatible with itself.
Remark 2.41.
  1. (i)
    If the domain of $\phi$ and $\psi$ is classical, then $\phi$ and $\psi$ are immediately compatible because the copy morphism is invariant under swapping:
[…]
  2. (ii)
    Compatibility of morphisms is clearly a symmetric relation. Indeed, if we assume $(\phi,\psi)=\operatorname{swap}\mathchoice{\,}{\,}{}{}(\psi,\phi)$, then we have
    $(\psi,\phi)=\operatorname{swap}\mathchoice{\,}{\,}{}{}\operatorname{swap}\mathchoice{\,}{\,}{}{}(\psi,\phi)=\operatorname{swap}\mathchoice{\,}{\,}{}{}(\phi,\psi),$
    and hence also $\psi$ and $\phi$ are compatible.
  3. (iii)
    Whenever $\phi$ and $\psi$ are self-adjoint, then compatibility is equivalent to $(\phi,\psi)$ being self-adjoint.
Example 2.42.
In $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$, two morphisms $\phi\colon A\to B$ and $\psi:A\to C$ are compatible if and only if
$\phi^{\mathrm{op}}(x)\,\psi^{\mathrm{op}}(y)=\psi^{\mathrm{op}}(y)\,\phi^{\mathrm{op}}(x)$
for all $x\in{}_{p}{B}$ and $y\in{}_{p}{C}$, which means that $\phi^{\mathrm{op}}$ and $\psi^{\mathrm{op}}$ have commuting ranges. In particular, $\phi$ is autocompatible if and only if the range of $\phi^{\mathrm{op}}$ is commutative. We can now explain the term ‘‘compatible’’ by showing how the of compatibility of quantum observables is an instance of this notion. If we take $B=C=\mathbb{C}^{\{0,1\}}$, then a self-adjoint morphism $\phi:A\to\mathbb{C}^{\{0,1\}}$ corresponds to a self-adjoint element $x\in{}_{p}{A}$ by Example 2.29, and similarly $\psi:A\to\mathbb{C}^{\{0,1\}}$ corresponds to a self-adjoint element $y\in{}_{p}{A}$. Then $\phi$ and $\psi$ are compatible if and only if the elements $x$ and $y$ commute, and this is the usual notion of compatibility of observables in quantum theory.
Another important aspect of compatibility is that it allows us to state a version of the universal property of the maximal tensor product of C*-algebras in all involutive Markov categories.
Proposition 2.43.
In any involutive Markov category, there is a bijective correspondence between the following two sets:
{Compatible pairs of deterministic $\mathopen{}\mathclose{{}\left\{\begin{array}[]{c}\text{Compatible pairs of deterministic morphisms}\\ A\to B\text{ and }A\to C\end{array}}\right\}\quad\stackrel{{\scriptstyle\cong}}{{\longleftrightarrow}}\quad\mathopen{}\mathclose{{}\left\{\begin{array}[]{c}\text{Deterministic morphisms}\\ A\to B\otimes C\end{array}}\right\}$ (2.10)
where a pair $(\phi\colon A\to B,\,\psi\colon A\to C)$ is sent to $(\phi,\psi)$.
In Proposition 3.12, we will turn to a slightly different version of this result, which actually reproduces the universal property of the maximal tensor product of C*-algebras. The present one does not, since here the morphisms are not required to be bounded in any way.
Proof.
We first prove that the map is well-defined by showing that such $(\phi,\psi)$ is indeed deterministic. Given a pair of compatible deterministic morphisms $\phi\colon A\to B$ and $\psi\colon B\to C$, by definition of compatibility $(\phi,\psi)$ is immediately self-adjoint. It is also deterministic, since the copy morphisms respect the symmetric monoidal structure (2.3).
Injectivity of the map (2.10) now immediately follows from marginalization, since $\phi$ and $\psi$ can be reconstructed from $(\phi,\psi)$ by composing with deletion on either output. For surjectivity, consider a deterministic morphism $\pi\colon A\to B\otimes C$, and take its marginals
[…]
As composites of deterministic morphisms, $\phi$ and $\psi$ are deterministic as well. Further we have
[…]
by determinism of $\pi$. Finally since $\pi$ is self-adjoint, $\phi$ and $\psi$ are also compatible by Remark 2.41(iii). ∎
Proposition 2.44.
Let $\phi,\psi$, and $\omega$ be pairwise compatible morphisms. Then $\phi$ is also compatible with $(\psi,\omega)$.
Proof.
This follows by associativity of the copy morphisms via the following direct calculation:
[…]
Corollary 2.45.
Let $\phi$ be an autocompatible morphism. Then $\phi^{(n)}$ is compatible with $\phi^{(m)}$ for any $n,m\in\mathbb{N}$.
Proof.
This follows by a double induction on $n$ and $m$ using Proposition 2.44. ∎
Autocompatibility has a strong relation with classicality, as the following results suggest.
Lemma 2.46.
Let $\phi\colon A\to B$ be an autocompatible morphism such that $\phi\otimes\phi$ is monic. Then $A$ is classical.
Proof.
This is a direct check:
[…]
The monicness assumption can now be applied to obtain the desired conclusion. ∎
Remark 2.47.
The mere existence of an autocompatible isomorphism $A\to B$ does not suffice to show that $B$ is classical. For instance in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ or $\operatorname{C*-}{\mathsf{pr}}_{\max}$, consider $B=M_{n}$, the matrix algebra of $n\times n$ matrices, and $A=\mathbb{C}^{n^{2}}$. Then there exist linear unital isomorphisms $\phi^{\mathrm{op}}\colon B\rightsquigarrow A$, and any such $\phi$ is autocompatible since it has commutative range.
Lemma 2.48.
Let $\phi\colon A\to B$ be a deterministic autocompatible epimorphism. Then $B$ is classical.
Proof.
By autocompatibility, $(\phi,\phi)$ is invariant under swap. Since $\phi$ is deterministic, we infer that
[…]
The epicness of $\phi$ implies that the same holds when $\phi$ is dropped; hence $B$ is classical. ∎
Combining Lemmas 2.46 and 2.48, we also obtain the following.
Corollary 2.49.
The domain and the codomain of a deterministic autocompatible isomorphism are classical.
Definition 2.50.
A morphism $\phi\colon A\to B$ in an involutive Markov category ${\mathsf{C}}$ is said to be non-invasive if it is compatible with the identity of $A$, i.e. if
$(\phi,\mathrm{id})=\operatorname{swap}\mathchoice{\,}{\,}{}{}(\mathrm{id},\phi).$
Remark 2.51.
From Remark 2.41 we deduce the following facts:
  1. (i)
    Any morphism with classical domain is non-invasive.
  2. (ii)
    Whenever $\phi$ is non-invasive and self-adjoint, then also $(\phi,\mathrm{id})$ is self-adjoint. In particular, the non-invasiveness of the identity morphism $\mathrm{id}_{A}$ of $A$ is equivalent to the classicality of the object $A$.
Moreover, a non-invasive morphism $\phi\colon A\to B$ is compatible with all morphisms out of $A$, as a simple string diagram calculation shows. In particular, every non-invasive morphism is also autocompatible.
Example 2.52.
A morphism $\phi\colon A\to B$ in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ or $\operatorname{C*-}{\mathsf{pr}}_{\max}$ is non-invasive if and only if
$\phi^{\mathrm{op}}(x)\,y=y\,\phi^{\mathrm{op}}(x)$
for all $x\in{}_{p}{B}$ and $y\in{}_{p}{A}$. In other words, we can equivalently characterize non-invasive morphisms as those that map to the center of ${A}$.
This notion of non-invasiveness is not to be confused with the concept of interaction-free measurement in quantum theory [46]. While a general definition of what counts as ‘‘interaction-free’’ does not seem to exist, we can see as follows that our notion of non-invasiveness is quite different. If $A$ is a matrix algebra (of matrix size $>1$), then its centre is trivial, and therefore the only non-invasive morphisms $A\to B$ are those that factor across $I=\mathbb{C}$. This is a physically uninteresting operation, as it amounts to discarding the original system and preparing a new state. But the case of matrix algebras is the one that most quantum information literature focuses on, including the literature on interaction-free measurements, which can still be interesting and nontrivial despite trivial center.
On the other hand, our notion of non-invasiveness also applies to morphisms that are not measurements. [^7] Measurements can be defined as morphisms $A\to B$ for which $B$ is classical; for $B=\mathbb{C}^{n}$ in $\mathsf{CPU}_{\min/\max}$ (see Definition 3.8), this reproduces the standard notion of POVM with finitely many outcomes.
Proposition 2.53.
Let $\phi$ and $\psi$ be any two morphisms in an involutive Markov category.
  1. (i)
    If $\phi$ and $\psi$ are non-invasive, then $\phi\otimes\psi$ is non-invasive.
  2. (ii)
    If $\phi$ and $\psi$ have the same domain, then both are non-invasive if and only if $(\phi,\psi)$ is non-invasive.
  3. (iii)
    If $\phi$ and $\psi$ are composable and $\phi$ is non-invasive, then $\psi\mathchoice{\,}{\,}{}{}\phi$ is non-invasive as well.
Proof.
Item (i) is proved using naturality of the swap and the monoidal multiplicativity of copy from (2.3). The ‘‘if’’ implication of Item (ii) follow by marginalization, while the ‘‘only if’’ is a consequence of Proposition 2.44. Concerning Item (iii), we have the following direct calculation:
{${\begin{tikzpicture}}\qed$
2.7 Almost sure equalities
Almost sure equality has been introduced in categorical probability in a special case in [7, Definition 5.1] and in [13, Definition 13.1] in general. We now develop this notion for involutive Markov categories based on Parzygnat’s [38, Definition 5.1], where the first two of the following variants were considered.
Definition 2.54.
Let $\omega\colon A\to B$ and $\phi,\psi\colon B\to C$ be morphisms in an involutive Markov category.
  1. (i)
[…]
  2. (ii)
[…]
  3. (iii)
[…]
Remark 2.55.
We emphasize that $\phi\cong_{\omega}\psi$ implies $\phi\simeq_{\omega}\psi$, simply by marginalizing the left and the right output of (2.12) separately. The converse is not true in general, as we will see in Example 3.20 as part of a more detailed study of almost sure equalities in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$.
Remark 2.56.
For $\omega\colon A\to B$ with $B$ classical, all four notions of almost sure equality coincide, as a direct consequence of $\operatorname{swap}_{B,B}\mathchoice{\,}{\,}{}{}\operatorname{copy}_{B}=\operatorname{copy}_{B}$.
We now recall the following tool that will be used in Section 3.2, and which has appeared as [38, Corollary 5.6].
Lemma 2.57.
If $\omega$, $\phi$ and $\psi$ are self-adjoint, then
[…]
In particular, whenever one of the two a.s. equalities holds, then $\phi\simeq_{\omega}\psi$ also holds.
Nevertheless, even in that case the symmetric a.s. equality may still be different (Example 3.20).
Lemma 2.58.
[…]
Proof.
[…]
3 Pictures, representability and the quantum de Finetti theorem
Involutive Markov categories like $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ contain ‘‘too many’’ morphisms to be directly relevant to quantum probability. This is already indicated by the isomorphisms $M_{n}\cong\mathbb{C}^{n^{2}}$ from Remark 2.47, but illustrated more drastically by the well-known no–broadcasting theorem [2]. In our setting, this states that the copy morphisms do not represent physically realizable processes. [^8] See semicartesian categorical probability [15, 24] for a different categorical approach to quantum probability which considers only physically realizable processes from the start. So involutive Markov categories constitute an abstract framework where we can broadcast, but this operation does not correspond to any physically implementable operation. [^9] See also [8], where the phrase logical broadcasting has been coined for this idea. For this reason, an involutive Markov category acts as an environment that enables string diagrammatic calculus, while physically realizable operations form a subcategory. This is analogous to how operator algebra provides an environment for quantum probability, while the physical observables are only the self-adjoint ones, and the physical operations are the completely positive ones. With these considerations in mind, in this section we introduce the notion of picture, which restricts from the environment to a subclass of morphisms, to be thought of as the physically realizable ones.
In contrast to involutive Markov categories, a picture is a good candidate for satisfying information flow axioms such as positivity and causality, as noted similarly by Parzygnat [38, Theorem 4.2 and Proposition 8.34]. In the following, we introduce and investigate a different kind of property, namely representability. In the classical case, this was defined in [17] as a way of describing a Markov category as the Kleisli category of an associated monad on its deterministic subcategory. For involutive Markov categories, we will argue that representability comes in two distinct flavors. The more straightforward one is given by extending the classical notion from [17, Definitions 3.7 and 3.10] to the involutive world. In the pictures that are of interest to quantum probability, the distribution objects (the objects defined by the universal property of representability) are then given by the universal C*-algebras [32] of an operator system, namely of the underlying operator system of the C*-algebra that we start with (see Proposition 3.36). The second flavour of representability is to require it only against classical objects. We call this concept classical representability. It describes the (C*-algebra of complex-valued continuous maps on the) state space of a C*-algebra through the associated universal property (Proposition 3.47). For this reason, we call state space objects the objects obtained from classical representability. A first interesting consequence of our synthetic approach is that the state space is Gelfand dual to the abelianization of the universal C*-algebra (Corollary 3.50).
Our final topic is the quantum de Finetti theorem. This result, which we prove for both the maximal and the minimal tensor products of C*-algebras in Appendix C, is central to our topic because of its inherent connection to classical representability, as it is categorically formulated as a factorization of exchangeable morphisms through the state space object (Theorem 3.64). This formulation is rather strong in that it not only applies to exchangeable states, but to exchangeable morphisms with general input and with an extra output as an additional tensor factor on the codomain.
3.1 Definition and main examples of pictures
Before introducing the concept of picture, let us consider a toy example that exists in every involutive Markov category.
Example 3.1 (The self-adjoint subcategory).
Let ${\mathsf{C}}$ be an involutive Markov category. We define $\operatorname{SA}({\mathsf{C}})$ to be the subcategory of self-adjoint morphisms. This subcategory has the following properties:
  1. (i)
    It is a symmetric monoidal subcategory containing the delete morphisms (by definition).
  2. (ii)
    Every morphism in it is self-adjoint (trivially).
  3. (iii)
    Whenever an autocompatible self-adjoint morphism $\phi\colon A\to B$ is compatible with a self-adjoint morphism $\psi\colon A\to C$, then also $(\phi,\psi)$ is self-adjoint, as a special case of Remark 2.41(iii).
These properties are the ones we now impose in the general definition of ‘‘picture’’.
Definition 3.2 (Picture).
Given an involutive Markov category ${\mathsf{C}}$, a wide subcategory [^10] This means that every object of ${\mathsf{C}}$ also belongs to ${\mathsf{D}}$, and implements the idea that a picture is a choice of morphisms rather than a choice of objects. ${\mathsf{D}}\subseteq{\mathsf{C}}$ is a picture in ${\mathsf{C}}$ if:
  1. (i)
    It is a symmetric monoidal subcategory containing the delete morphisms.
  2. (ii)
    Every morphism of ${\mathsf{D}}$ is self-adjoint.
  3. (iii)
    Whenever an autocompatible morphism $\phi\colon A\to B$ (of ${\mathsf{D}}$) is compatible with a morphism $\psi\colon A\to C$ (of ${\mathsf{D}}$), then also $(\phi,\psi)$ belongs to ${\mathsf{D}}$.
The last property is a weakening of the property discussed in Remark 2.41(iii). This weakening is relevant since if we used the stronger formulation of Remark 2.41(iii) in the definition, then we would exclude the minimal C*-tensor norm as a possible monoidal structure (Proposition 3.12).
When dealing with pictures, we reserve the term morphism for the morphisms in ${\mathsf{D}}$, while morphisms in ${\mathsf{C}}$ will be called generalized morphisms. Generalized morphisms in string diagrams will be denoted by dashed boxes:
[…]
The copy morphism form an exception: although these typically do not belong to the picture, we will still use the same notation as before.
Further, the term state will refer to the morphisms (not generalized!) whose domain is the monoidal unit. In string diagrams, this type of morphism is depicted as
[…]
to emphasize that such morphisms ‘‘have no input’’.
For brevity, we usually leave ${\mathsf{C}}$ implicit and say that ${\mathsf{D}}$ is a picture. Whenever the containing category ${\mathsf{C}}$ needs to be referenced, we will denote it by $\operatorname{Gen}({\mathsf{D}})$. This reminds us of the idea that $\operatorname{Gen}({\mathsf{D}})$ is the category of generalized morphisms of ${\mathsf{D}}$. Also, we will use the shorthand
${\mathsf{D}}_{\mathrm{det}}\coloneqq{\mathsf{D}}\cap{\mathsf{C}}_{\mathrm{det}}$
to indicate the symmetric monoidal subcategory of deterministic morphisms.
Remark 3.3.
As suggested to us by Sam Staton, a picture could alternatively be defined in terms of a strong symmetric monoidal faithful functor $U\colon{\mathsf{D}}\to\operatorname{SA}({\mathsf{C}})$ rather than in terms of ${\mathsf{D}}$ being a subcategory. This version may be preferred for better analogy with forgetful functors. However, we have not adopted this definition as it would clutter our notation further, while being essentially equivalent: $U$ can always be taken to be such an inclusion functor, at least modulo replacing $\operatorname{Gen}({\mathsf{D}})$ by an equivalent involutive Markov category.
Definition 3.4.
A picture ${\mathsf{D}}$ displays compatibility if, whenever $\phi$ and $\psi$ are compatible, then $(\phi,\psi)$ is a morphism in ${\mathsf{D}}$.
Remark 3.5.
Property (iii) of Definition 3.2 already allows some interesting ‘‘displaying’’. For instance, if a morphism $\phi$ is non-invasive, then $(\phi,\mathrm{id})$ is also a morphism by using the mentioned property with $\psi=\mathrm{id}$.
Another example is given by autocompatibility: if $\phi$ is an autocompatible morphism, an inductive argument combining Corollary 2.45 and Property (iii) shows that $\phi^{(n)}$, for any positive integer $n$, also belongs to the picture. In fact, this can also be reversed using Remark 2.41(iii): if $\phi^{(n)}$ is a morphism for some $n\geq 2$, then $\phi$ is autocompatible.
Remark 3.6.
The classical subcategory of a picture ${\mathsf{D}}$, defined as
$C\ell({\mathsf{D}})\coloneqq{\mathsf{D}}\cap C\ell(\operatorname{Gen}({\mathsf{D}})),$
is a Markov category. Indeed, by Property (iii) of Definition 3.2, Remark 2.41(i) implies that the copy morphisms belong to $C\ell({\mathsf{D}})$.
Remark 3.7.
Given any involutive Markov category ${\mathsf{C}}$, the self-adjoint subcategory $\operatorname{SA}({\mathsf{C}})$ is a picture that displays compatibility: see Remark 2.41(iii).
However, in our motivating examples, the pictures of interest are a proper subcategory of the self-adjoint morphisms.
Definition 3.8.
In the category $\operatorname{C*-}{\mathsf{pr}}$, we denote by $\mathsf{CPU}$ the subcategory whose morphisms are (formal opposites of) completely positive unital linear maps.
For the definition of such maps, we refer to Definition A.8. Whenever $\mathsf{CPU}$ is considered as a subcategory of either $\operatorname{C*-}{\mathsf{pr}}_{\min}$ or $\operatorname{C*-}{\mathsf{pr}}_{\max}$, we will decorate it accordingly: $\mathsf{CPU}_{\min}$ and $\mathsf{CPU}_{\max}$, respectively. For brevity, we also use $\mathsf{CPU}_{\min/\max}$ to make statements that apply to both cases.
As anticipated in Remark 2.32, $\mathsf{CPU}_{\min}$ and $\mathsf{CPU}_{\max}$ are different categories, since completely positive maps are bounded, and therefore influenced by the choice of the norm. This difference manifests itself also in formal categorical properties in Proposition 3.12.
Notation 3.9.
From now on, for cleaner writing we sometimes prefer working with general elements of the containing C*-algebras $A$ rather than with elements of the dense ***-subalgebras ${}_{p}{A}$. We employ this only in contexts where all maps involved are positive unital, since such maps have a unique extension to the containing C*-algebras (see Lemma A.4).
Remark 3.10.
In [38], the focus is on subcategories that are not necessarily symmetric monoidal. In fact, the information flow axioms considered there are obtained by the Kadison–Schwarz inequality (Corollary A.11), and all results can therefore be stated in the context of positive unital maps satisfying such an inequality. However, there does not seem to be any concrete advantage in considering such generality, since the physically realizable operations are only the completely positive unital maps, and also the mathematical literature mostly focuses on these. Furthermore, considering only symmetric monoidal subcategories provides a more understandable playground, as the reader does not have to pay attention to when tensoring is allowed. An additional motivation is given by the proof of Proposition 3.11: Property (iii) holds provided that the morphisms considered are completely positive, since without this hypothesis the universal property of the maximal tensor product [5, Exercise 3.5.1] could not be applied.
Proposition 3.11.
The subcategories $\mathsf{CPU}_{\min/\max}$ are pictures.
Proof.
All morphisms are self-adjoint by Proposition A.9. Moreover, the subcategories $\mathsf{CPU}_{\min/\max}$ are symmetric monoidal since the coherence morphisms are bounded [5, Exercises 3.1.2, 3.3.1 and 3.3.2] (recall Lemma A.4), and the tensor product is respected by Proposition A.12. Moreover, the unique unital map $\mathbb{C}\rightsquigarrow{}_{p}{A}$ corresponding to deletion is completely positive.
Concerning Property (iii), we recall Example 2.42: For any autocompatible morphism $\phi\colon A\to B$, the completely positive map $\phi^{\mathrm{op}}$ has commutative range. In particular, the range $\operatorname{\operatorname{im}}(\phi^{\mathrm{op}})$ generates a commutative C*-subalgebra $R\subseteq A$, and so $\phi$ admits a factorization $\phi=\phi^{\prime}\mathchoice{\,}{\,}{}{}\iota$, where $\iota^{\mathrm{op}}\colon R\rightsquigarrow A$ is the inclusion. Since any morphism ${\psi}\colon A\to C$ compatible with $\phi$ is also compatible with $\iota$, the map $(\iota,\psi)^{\mathrm{op}}$ is therefore completely positive because commutative C*-algebras are nuclear [40, Proposition 12.9], so that we can apply the universal property of the maximal tensor product [5, Exercise 3.5.1] for both monoidal structures. We conclude that $(\phi,\psi)$ is the composition of two morphisms $\phi^{\prime}\otimes\mathrm{id}$ and $(\iota,\psi)$ which are both completely positive, and so $(\phi,\psi)$ is completely positive as well. ∎
Proposition 3.12.
$\mathsf{CPU}_{\max}$ displays compatibility, while $\mathsf{CPU}_{\min}$ does not.
Proof.
Since compatibility is equivalent to commuting ranges in $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$ (Example 2.42), this is the universal property of the maximal tensor product: [5, Exercises 3.5.1 and 3.6.3] explain why $\mathsf{CPU}_{\max}$ has this property and $\mathsf{CPU}_{\min}$ does not. ∎
One may wonder why we have not required ${\mathsf{D}}_{\mathrm{det}}=\operatorname{Gen}({\mathsf{D}})_{\mathrm{det}}$ in the definition of picture. Indeed, between C*-algebras all ***-homomorphisms are bounded. However, this is no longer true for C*-pairs, because we are missing the completeness of the normed spaces. A simple example is the following.
Example 3.13 (Boundedness is not automatic for ***-homomorphisms).
It is well-known that a positive unital map is always bounded (Proposition A.9). However, the same is not true for ***-homomorphisms. Equivalently, not every ***-homomorphism between pre-C*-algebras is positive.
For instance, consider two C*-pairs $A$ and $B$. Then there is a generalized morphism $\iota\colon A\otimes_{\max}B\to A\otimes_{\min}B$, where the subscripts indicate the norm we are considering, given by the identity ${}_{p}{A}\odot{}_{p}{B}\rightsquigarrow{}_{p}{A}\odot{}_{p}{B}$. We claim that this ***-homomorphism is not bounded in general.
Indeed, if $\iota^{\mathrm{op}}$ was bounded for all choices of ${}_{p}{A}$ and ${}_{p}{B}$, then we would get a ***-homomorphism $\operatorname{cl}({\iota^{\mathrm{op}}})\colon A\otimes_{\min}B\to A\otimes_{\max}B$ by Lemma A.4. This $\operatorname{cl}({\iota^{\mathrm{op}}})$ would be unital and positive by Proposition A.10, so that $\mathopen{}\mathclose{{}\left\lVert\iota^{\mathrm{op}}}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\operatorname{cl}({\iota^{\mathrm{op}}})}\right\rVert=1$. In particular, we would get that $\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\max}\leq\mathopen{}\mathclose{{}\left\lVert\sum_{i}a_{i}\otimes b_{i}}\right\rVert_{\min}$ for all elements of the algebraic tensor product. Since the reverse inequality always holds, this would mean that the two norms coincide, which is generally not the case [45, Theorem 6].
A more concrete example is the following. Let us consider $A=\mathbb{C}$ and $(B,{}_{p}{B})=({C}([0,1]),\mathbb{C}[x])$, where $C([0,1])$ is the C*-algebra of complex-valued continuous maps on the unit interval $[0,1]$, and $\mathbb{C}[x]$ is included in $C([0,1])$ by identifying each polynomial $f\in\mathbb{C}[x]$ with the associated function $x\mapsto f(x)$. By the Stone–Weierstrass theorem, such an identification makes $\mathbb{C}[x]$ a dense ***-subalgebra of $C([0,1])$. Let us now take $\phi^{\mathrm{op}}\colon\mathbb{C}[x]\rightsquigarrow\mathbb{C}$ given by evaluating polynomials at any fixed real number outside of $[0,1]$. This is a ***-homomorphism, but it cannot be bounded, because by Gelfand duality every bounded ***-homomorphism $C([0,1])\rightsquigarrow\mathbb{C}$ is given by evaluation at a point of $[0,1]$.
3.2 Almost sure equalities via nullspaces
Almost sure equality is a central notion in probability, and we have already considered the various definitions in the involutive setting in Section 2.7. We devote this subsection to studying the meaning of these almost sure equalities in $\mathsf{CPU}_{\min/\max}$.
Definition 3.14.
Let $\omega^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ be a completely positive unital map between pre-C*-algebras. Then:
  1. (i)
    Its left nullspace (resp. right nullspace) is the subset of ${}_{p}{B}$ given by
[…]
  2. (ii)
    Its symmetric nullspace is the subset of ${}_{p}{B}$ given by
[…]
Lemma 3.15.
The left (resp. right) nullspace is the largest closed right (resp. left) ideal contained in the kernel of $\omega^{\mathrm{op}}$, and
$x\in N_{\omega^{\mathrm{op}}}\qquad\Longleftrightarrow\qquad x^{*}\in{}_{\omega^{\mathrm{op}}}N.$
Proof.
First, ${}_{\omega^{\mathrm{op}}}N$ (resp. $N_{\omega^{\mathrm{op}}}$) is a right (resp. left) ideal by the inequality $0\leq yy^{*}\leq\mathopen{}\mathclose{{}\left\lVert y}\right\rVert^{2}$ (Corollary A.18) and positivity of $\omega^{\mathrm{op}}$. The closedness is straightforward. To prove that both ideals are contained in the kernel of $\omega^{\mathrm{op}}$, we note that, by the Kadison–Schwarz inequality (Corollary A.11),
$0=\omega^{\mathrm{op}}(xx^{*})\geq\omega^{\mathrm{op}}(x)\,\omega^{\mathrm{op}}(x)^{*}\geq 0$
Hence these inequalities must be equalities, and $\omega^{\mathrm{op}}(x)=0$. For the maximality, let $J$ be a right ideal contained in the kernel of $\omega^{\mathrm{op}}$. Then we indeed must have $J\subseteq{}_{\omega^{\mathrm{op}}}N$, because $x\in J$ gives $xx^{*}\in J$, and therefore $\omega^{\mathrm{op}}(xx^{*})=0$. The final statement holds by definition. ∎
Lemma 3.16.
The symmetric nullspace $SN_{\omega^{\mathrm{op}}}$ is the largest closed two-sided ideal contained in the kernel of $\omega^{\mathrm{op}}$, and it coincides with
[…]
Moreover, it is a ***-ideal, i.e. $x\in SN_{\omega^{\mathrm{op}}}$ if and only if $x^{*}\in SN_{\omega^{\mathrm{op}}}$.
Proof.
$SN_{\omega^{\mathrm{op}}}$ is a two-sided ideal by the following general fact applied to $J=N_{\omega^{\mathrm{op}}}$: Whenever $J$ is a left ideal in a ring $R$, then the ideal quotient
[…]
is clearly a two-sided ideal. It is closed by construction as the intersection of the closed sets
[…]
Moreover, it is contained in the kernel of $\omega^{\mathrm{op}}$ because $SN_{\omega^{\mathrm{op}}}\subseteq N_{\omega^{\mathrm{op}}}$, and it is the largest two-sided ideal contained in the kernel since any other such ideal $J$ must satisfy $J\subseteq N_{\omega^{\mathrm{op}}}$ by Lemma 3.15 and therefore $J\subseteq SN_{\omega^{\mathrm{op}}}$ by the closedness under multiplication from the right. This maximality property also implies that $SN_{\omega^{\mathrm{op}}}$ coincides with the given $S{}_{\omega^{\mathrm{op}}}N$. The final statement follows from $SN_{\omega^{\mathrm{op}}}=S{}_{\omega^{\mathrm{op}}}N$. ∎
Lemma 3.17.
If $\omega^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ is a completely positive unital map and $\omega^{\mathrm{op}}(x^{*}x)=0$ for some $x\in{}_{p}{B}$, then also
$\omega^{\mathrm{op}}(x^{*}y)=0=\omega^{\mathrm{op}}(y^{*}x)$
for every $y\in{}_{p}{B}$.
Proof.
By Lemma 3.15, the right nullspace is a left ideal contained in the kernel of $\omega^{\mathrm{op}}$. This means that $x\in N_{\omega^{\mathrm{op}}}$ implies $y^{*}x\in N_{\omega^{\mathrm{op}}}$, and in particular $\omega^{\mathrm{op}}(y^{*}x)=0$. Self-adjointness of $\omega^{\mathrm{op}}$ now implies $\omega^{\mathrm{op}}(x^{*}y)=0$ as well. ∎
The nullspaces now give us the following characterization of the almost sure equalities, extending results of Parzygnat [38, Theorem 5.12].
Theorem 3.18.
Let us consider a morphism $\omega\colon A\to B$ in $\mathsf{CPU}_{\min/\max}$ and two generalized morphisms $\phi,\psi\colon B\to C$. Then:
[…]
Proof.
The first two characterizations are already covered in [38, Theorem 5.12]. [^11] Although that reference is concerned with finite-dimensional C*-algebras only, the same proof works in general. The third one uses a similar argument as follows. We first note that $\phi\cong_{\omega}\psi$ can be spelled out explicitly as
$\omega^{\mathrm{op}}(z_{1}\phi^{\mathrm{op}}(y)z_{2})=\omega^{\mathrm{op}}(z_{1}\psi^{\mathrm{op}}(y)z_{2})$
for all $z_{1},z_{2}\in{}_{p}{B}$ and $y\in{}_{p}{C}$, or equivalently $\omega^{\mathrm{op}}(z_{1}(\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y))z_{2})=0$. By choosing
$z_{1}\coloneqq z_{2}^{*}(\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y))^{*},$
we conclude $\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y)\in SN_{\omega^{\mathrm{op}}}$. Conversely, if we set $x\coloneqq\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y)\in SN_{\omega^{\mathrm{op}}}$, then $\omega^{\mathrm{op}}((xz_{2})^{*}xz_{2})=0$ for every $z_{2}\in{}_{p}{B}$, and Lemma 3.17 implies that $\omega^{\mathrm{op}}(z_{1}xz_{2})=0$ for any choice of $z_{1}\in{}_{p}{B}$. ∎
Remark 3.19.
Consider a simple pre-C*-algebra ${}_{p}{B}$, i.e. such that the only closed two-sided ideals are $\{0\}$ and ${}_{p}{B}$ itself. Then for every completely positive unital map $\omega^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$, the symmetric nullspace $SN_{\omega^{\mathrm{op}}}$ must be trivial, since it is a closed two-sided ideal not containing $1$. By Theorem 3.18, for every generalized $\phi,\psi\colon B\to C$, it follows that
$\phi\cong_{\omega}\psi\quad\Longleftrightarrow\quad\phi=\psi$
whenever ${}_{p}{B}$ is simple.
Example 3.20 (Almost sure equalities are not necessarily symmetric).
In general, $\phi\simeq_{\omega}\psi$ does not imply $\phi\cong_{\omega}\psi$. The following counterexample is a special case of [39, Remark 2.75]. For the sake of simplicity, we just consider the case of the morphisms defined below, where all omitted matrix entries are zero.
$\begin{array}[]{rclcrclcrcl}\omega^{\mathrm{op}}\colon M_{4}&\rightsquigarrow&\mathbb{C}&&\phi^{\mathrm{op}}\colon M_{2}&\rightsquigarrow&M_{4}&&\psi^{\mathrm{op}}\colon M_{2}&\rightsquigarrow&M_{4}\\ x&\mapsto&\operatorname{tr}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\begin{smallmatrix}1&&&\\ &1&&\\ &&0&\\ &&&0\end{smallmatrix}}\right)\cdot x}\right)&&x&\mapsto&\mathopen{}\mathclose{{}\left(\begin{smallmatrix}x&&\\ &\operatorname{tr}(x)&\\ &&\operatorname{tr}(x)\end{smallmatrix}}\right)&&x&\mapsto&\mathopen{}\mathclose{{}\left(\begin{smallmatrix}x&\\ &x\end{smallmatrix}}\right)\end{array}$
[…]
Lemma 3.21.
Let us consider $\mathsf{CPU}_{\min/\max}$ and a morphism $\omega\colon A\to B$. Then the following are equivalent:
  1. (i)
[…]
  2. (ii)
    $N_{\omega^{\mathrm{op}}}={}_{\omega^{\mathrm{op}}}N$.
Proof.
The implication (ii) $\Rightarrow$ (i) is an immediate application of Theorem 3.18, so we are left to show the converse. For the other direction, by Theorem 3.18 it is enough to show that every $x\in{}_{p}{A}$ can be written as $x=\phi^{\mathrm{op}}(y)-\psi^{\mathrm{op}}(y)$ for some $y\in{}_{p}{C}$ for suitable generalized morphisms $\phi,\psi\colon B\to C$ and suitable $C$. For this purpose, we take $C=\mathbb{C}^{\{0,1\}}$ and $y=e_{1}\coloneqq(0,1)$. Taking $C=\mathbb{C}^{\{0,1\}}$ and $y=e_{1}\coloneqq(0,1)$, this is clear from Example 2.29, which lets us achieve $\phi^{\mathrm{op}}(e_{1})=0$ and $\psi^{\mathrm{op}}(e_{1})=x$. ∎
Remark 3.22.
In the proof of Lemma 3.21, $\phi$ and $\psi$ are generally not self-adjoint because of Lemma 2.57.
Remark 3.23.
By looking at Lemma 3.21 and Lemma 2.58, one may wonder whether $N_{\omega^{\mathrm{op}}}={}_{\omega^{\mathrm{op}}}N$ holds if and only if $N_{\omega^{\mathrm{op}}}=SN_{\omega^{\mathrm{op}}}$. This is indeed the case, because $N_{\omega^{\mathrm{op}}}={}_{\omega^{\mathrm{op}}}N$ implies that $N_{\omega^{\mathrm{op}}}$ is a closed two-sided ideal, and hence we have $N_{\omega^{\mathrm{op}}}\subseteq SN_{\omega^{\mathrm{op}}}$ by the maximality statement of Lemma 3.16. On the other hand, $SN_{\omega^{\mathrm{op}}}\subseteq N_{\omega^{\mathrm{op}}}$ holds trivially.
3.3 Kolmogorov products
In classical Markov categories, the notion of Kolmogorov product axiomatizes the idea of taking infinite tensor products of objects [20]. This is relevant when talking about joint distributions of infinitely many random variables. We now generalize this notion to the involutive setting in terms of pictures.
Definition 3.24.
Let ${\mathsf{D}}$ be a picture. For any set $J$, a Kolmogorov product of a family of objects $(A_{j})_{j\in J}$ in ${\mathsf{D}}$ is an object $A_{J}$ together with a natural bijection between
  1. (i)
    generalized morphisms $B\to A_{J}\otimes E$, and
  2. (ii)
    families of generalized morphisms $(\phi_{F}\colon B\to A_{F}\otimes E)$, where $A_{F}$ is shorthand for $\bigotimes_{j\in F}A_{j}$, indexed by finite subsets $F\subseteq J$, such that whenever $F^{\prime}\subseteq F$, we have
[…]
    where the projection $\pi_{F,F^{\prime}}\colon A_{F}\to A_{F^{\prime}}$ is the identity on $A_{F^{\prime}}$ and the delete morphism in the other components.
Moreover, this natural bijection is required to restrict to natural bijections on ${\mathsf{D}}$ and $\operatorname{Gen}({\mathsf{D}})_{\mathrm{det}}$.
As the definition is in terms of a universal property, the Kolmogorov product is unique up to a unique isomorphism in ${\mathsf{D}}$. The Kolmogorov product of a constant family $(A)_{j\in J}$ is called Kolmogorov power.
Definition 3.25.
A picture ${\mathsf{D}}$ has (countable) Kolmogorov products if every (countable) family of objects has a Kolmogorov product.
The notion of Kolmogorov products has already appeared implicitly in the study of the quantum de Finetti theorem by Staton and Summers [42, Definition 2.23 and Theorem 2.24].
Remark 3.26.
Let us consider a picture with countable Kolmogorov products and a morphism $\phi\colon A\to B$. Then we can construct a generalized morphism $\phi^{(\mathbb{N})}\colon A\to B_{\mathbb{N}}$ obtained by the family $(\phi^{(n)})$, where $n$ is any nonnegative integer.
Then $\phi^{(\mathbb{N})}$ is a morphism if and only if $\phi$ is autocompatible. This is immediate from Remark 3.5.
Proposition 3.27.
The pictures $\mathsf{CPU}_{\min/\max}$ have Kolmogorov products.
Proof.
We use the same notation as in Definition 3.24 above. If some $A_{j}$ is the zero algebra, then it is easy to see that the zero algebra also serves as a Kolmogorov product $A_{J}$. We can therefore assume that all $A_{j}$ are non-zero.
Then we need to consider the diagram formed by the finite tensor products $A_{F}\coloneqq\bigotimes_{j\in F}A_{j}$, for any finite subset $F\subseteq J$, and for any $F^{\prime}\subseteq F$, the canonical morphisms $\pi_{F,F^{\prime}}\colon A_{F}\to A_{F^{\prime}}$ corresponding to the bounded inclusion ***-homomorphisms
$\pi{}^{\mathrm{op}}_{F,F^{\prime}}\colon{}_{p}{A}_{F^{\prime}}\rightsquigarrow{}_{p}{A}_{F}.$
Our assumption that all $A_{j}$ are non-zero implies that these ***-homomorphisms are isometries. In $\operatorname{C*-}{\mathsf{pr}}$, the filtered colimit of this diagram can be constructed as
[…]
where, given $x_{F}\in{}_{p}{A_{F}}$ and $x_{F^{\prime}}\in{}_{p}{A_{F^{\prime}}}$, we have $x_{F}\sim x_{F^{\prime}}$ if and only if there exists $G\supseteq F,F^{\prime}$ such that $\pi{}^{\mathrm{op}}_{G,F}(x_{F})=\pi{}^{\mathrm{op}}_{G,F^{\prime}}(x_{F^{\prime}})$. With the induced algebraic structure, this clearly makes ${}_{p}{A_{J}}$ the colimit both in the category of ***-algebras and in the category of vector spaces. Since all $\pi{}_{F,F^{\prime}}$ are isometries, the colimit also inherits a C*-norm from the ${}_{p}{A_{F}}$. Its completion $A_{J}$ is the colimit of the $A_{F}$ in the category of C*-algebras, since this is exactly how filtered colimits of C*-algebras can be constructed [36, Section 6.1] [^12] See also the more general [6, Corollary 7.22], or the historical reference [23, Chapter 2]..
So the desired bijection holds with $E=\mathbb{C}$ both at the level of generalized morphisms and at the level of generalized deterministic morphisms. It is also easy to see that the colimit is preserved by $-\otimes E$ for any C*-pair $E$. By direct check, this bijection restricts to ***-homomorphisms, which are the deterministic generalized morphisms.
To show that we also obtain bijections at the level of pictures, let us consider a completely positive unital map ${}_{p}{A_{J}}\odot{}_{p}{E}\rightsquigarrow{}_{p}{B}$. Then the composite
${}_{p}{A_{F}}\odot{}_{p}{E}\rightsquigarrow{}_{p}{A_{J}}\odot{}_{p}{E}\rightsquigarrow{}_{p}{B}$
is a completely positive unital map for any $F$ because it is a composition of two such morphisms. Conversely, let us assume that all $\phi^{\mathrm{op}}_{F}\colon{}_{p}{A_{F}}\odot{}_{p}{E}\rightsquigarrow{}_{p}{B}$ are completely positive unital maps (for $F\subseteq J$ finite). Then the induced $\phi^{\mathrm{op}}\colon{}_{p}{A_{J}}\odot{}_{p}{E}\rightsquigarrow{}_{p}{B}$ has norm $\leq 1$: given any $[x_{F}]\in{}_{p}{A}_{J}\odot{}_{p}{E}$ such that $x_{F}\in{}_{p}{A}_{F}\odot{}_{p}{E}$,
$\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}([x_{F}])}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\phi_{F}^{\mathrm{op}}(x_{F})}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x_{F}}\right\rVert=\mathopen{}\mathclose{{}\left\lVert[x_{F}]}\right\rVert.$
By Proposition A.9, $\phi^{\mathrm{op}}$ is positive. Since $B$ and $E$ are arbitrary, also $\phi^{\mathrm{op}}\odot\mathrm{id}_{C}$ is positive for any $C$ because $\phi^{\mathrm{op}}_{F}\odot\mathrm{id}_{C}$ is. Therefore, $\phi^{\mathrm{op}}$ is completely positive. ∎
Notation 3.28.
Let $J$ be any set. We recall that a finite permutation is a permutation $J\to J$ that fixes all but finitely many elements.
Let us consider a picture ${\mathsf{D}}$, and suppose that the Kolmogorov power $A_{J}$ of an object $A$ exists. Then for every finite permutation $\sigma$, the universal property of $A_{J}$ induces a morphism
$A_{\sigma}\colon A_{J}\to A_{J}$
that swaps the occurrences of $A$ in $A_{J}$ according to $\sigma$.
Remark 3.29 (Zero-one laws).
Some readers may wonder whether the zero–one laws of [20] generalize to the involutive setting. Indeed the abstract Kolmogorov zero–one law [20, Theorem 5.3] holds with the same proof, as long as one takes care to preserve the order of the tensor factors throughout the argument, and where all morphisms in the statement and proof are allowed to be generalized morphisms. At the same time, this vast generality also indicates that Kolmogorov’s zero–one law is not a particularly deep or interesting result, which is why we do not elaborate further.
The abstract Hewitt–Savage zero–one law [20, Theorem 5.4] is a bit deeper, and Parzygnat already asked for a quantum generalization of it [38, Question 8.41]. Unfortunately, this zero–one law is essentially vacuously true even in the whole involutive Markov categories $\operatorname{C*-}{\mathsf{pr}}_{\min}$ and $\operatorname{C*-}{\mathsf{pr}}_{\max}$. The reason is that for infinite $J$, every generalized morphism $\phi\colon A_{J}\to B$ which satisfies $\phi\mathchoice{\,}{\,}{}{}A_{\sigma}=\phi$ for any finite permutation $\sigma$ of $J$ already factors through the delete morphism. Indeed for such $\phi$, every $\phi^{\mathrm{op}}(y)\in{}_{p}{A_{J}}$ for $y\in{}_{p}{B}$ must be invariant under finite permutations. Now the construction of ${}_{p}{A_{J}}$ implies that $\phi^{\mathrm{op}}(y)=[x]$ for some $x\in{}_{p}{A_{F}}$ and some finite subset $F\subseteq J$. By choosing any finite permutation $\sigma\colon J\to J$ such that $F\cap\sigma(F)=\varnothing$, we conclude that $x=\lambda 1$, so that $\phi$ in fact factors through the delete morphism.
One may wonder if such an unfortunate turn of events happens because we are working with C*-pairs instead of C*-algebras. This is not the case as the issue still persists for tensor powers of C*-algebras (Proposition B.1).
3.4 Representability
We start with the definition of representability. The reader may refer to [17] for representability of classical Markov categories; here we sketch most of the ideas and generalize them to the involutive setting.
Definition 3.30.
Let ${\mathsf{D}}$ be a picture. Given an object $A$, a distribution object is an object $PA$ together with a sampling morphism $\mathsf{samp}_{A}\colon PA\to A$ such that the induced map
$\mathsf{samp}_{A}\circ\_\colon{\mathsf{D}}_{\mathrm{det}}(B,PA)\to{\mathsf{D}}(B,A)$
is bijective for every object $B\in{\mathsf{D}}$. We say that ${\mathsf{D}}$ is representable if every object $A$ admits a distribution object $PA$.
Notation 3.31.
We will denote by $(\_)^{\sharp}$ the inverse of $\mathsf{samp}_{A}\circ\_$, and $\delta_{A}\coloneqq(\mathrm{id}_{A})^{\sharp}\colon A\to PA$ is also called the delta morphism. We will drop the subscript $A$ if it is sufficiently clear from context.
Our first observation extends [17, Lemma 3.9] to the involutive setting.
Lemma 3.32.
Let ${\mathsf{D}}$ be a representable picture. Then $A\mapsto PA$ extends to a functor ${\mathsf{D}}\to{\mathsf{D}}_{\mathrm{det}}$ which is right adjoint to the inclusion ${\mathsf{D}}_{\mathrm{det}}\hookrightarrow{\mathsf{D}}$ with:
  1. (i)
    Unit given by the delta morphisms;
  2. (ii)
    Counit given by the sampling morphisms.
Proof.
Representability is exactly the requirement of ${\mathsf{D}}_{\mathrm{det}}\hookrightarrow{\mathsf{D}}$ being left adjoint, using the definition via universal morphisms with $P$ as the right adjoint. ∎
Notation 3.33.
In view of Lemma 3.32, we write $P\phi$ to denote the morphism $PA\to PB$ obtained from $\phi\colon A\to B$ via the functor $P$.
Proposition 3.34.
Let ${\mathsf{D}}$ be a representable picture. Then the right adjoint $P\colon{\mathsf{D}}\to{\mathsf{D}}_{\mathrm{det}}$ has a canonical symmetric lax monoidal structure given by
[…]
and the adjunction between $P$ and the inclusion $\operatorname{incl}\colon{\mathsf{D}}_{\mathrm{det}}\to{\mathsf{D}}$ is symmetric monoidal.
Proof.
The requirement of ${\mathsf{D}}$ being monoidal makes sure that ${\mathsf{D}}_{\mathrm{det}}\hookrightarrow{\mathsf{D}}$ is a strict symmetric monoidal functor. The rest of the proof is exactly as in [17, Proposition 3.15]. ∎
From this result, we get a corollary analogous to [17, Corollary 3.17].
Corollary 3.35.
Let ${\mathsf{D}}$ be a representable picture. The adjunction given by $P\colon{\mathsf{D}}\to{\mathsf{D}}_{\mathrm{det}}$ and the inclusion $\operatorname{incl}\colon{\mathsf{D}}_{\mathrm{det}}\hookrightarrow{\mathsf{D}}$ induces a symmetric monoidal and affine [^13] A monad on a category with terminal object $I$ is affine if $P(I)\cong I$. monad $P\operatorname{incl}$ on ${\mathsf{D}}_{\mathrm{det}}$. In particular, ${\mathsf{D}}$ and the Kleisli category $K\ell(P\operatorname{incl})$ are isomorphic as symmetric monoidal categories.
We can now construct this monad for our two pictures involving completely positive maps.
Proposition 3.36.
The pictures $\mathsf{CPU}_{\min/\max}$ are representable.
Proof.
In the context of C*-algebras and completely positive unital maps, the universal property of $PA$ translates to that of universal C*-algebras, where the C*-algebra $A$ is seen as an operator system [32, Section 3]. Let us recall the construction of the universal C*-algebra described in [30, before Lemma 2.5]. We consider the free unital ***-algebra
$FA\coloneqq\bigoplus_{n\geq 0}{A^{\odot n}}$
with multiplication given by tensoring and star given by reversing the order of every elementary tensor $x_{1}\odot\cdots\odot x_{n}$, applying the star on $A$ to each element, and extending linearly. For every completely positive unital map $\phi^{\mathrm{op}}\colon A\rightsquigarrow B$, we now obtain a ***-homomorphism $\widetilde{\phi}^{\mathrm{op}}\colon F{A}\rightsquigarrow{B}$ by setting
$\widetilde{\phi}^{\mathrm{op}}(x_{1}\odot\cdots\odot x_{n})\coloneqq\phi^{\mathrm{op}}(x_{1})\cdots\phi^{\mathrm{op}}(x_{n})$
and extending linearly. This induces a seminorm on $FA$ given by
$\mathopen{}\mathclose{{}\left\lVert x}\right\rVert_{PA}\coloneqq\sup_{\phi}\,\mathopen{}\mathclose{{}\left\lVert\widetilde{\phi}^{\mathrm{op}}(x)}\right\rVert,$
where $\phi$ ranges over completely positive unital maps $A\rightsquigarrow B$ for any $B$. Finally, we take $PA$ to be the completion of $FA$, which is now a C*-algebra. The map $\phi^{\sharp,\mathrm{op}}\colon PA\rightsquigarrow B$ is then the one induced by $\widetilde{\phi}^{\mathrm{op}}$. The universal property of the universal C*-algebra is now encoded in the completely positive unital inclusion $A\rightsquigarrow PA$, in the sense that it is the initial completely positive unital map from $A$ to a C*-algebra.
To get a C*-pair $(PA,{}_{p}{PA})$, we define ${}_{p}{PA}$ as the dense ***-subalgebra of $PA$ generated by the image of ${}_{p}{A}$ via $A\rightsquigarrow PA$. By definition, ${}_{p}{PA}$ can equivalently be seen as a quotient of the free unital ***-algebra $F{}_{p}{A}$ associated to ${}_{p}{A}$ instead of $A$. Therefore, given any completely positive unital map ${}_{p}{A}\rightsquigarrow{}_{p}{B}$, its closure gives a completely positive unital map $A\rightsquigarrow B$ (Lemma A.4), and so the map ${}_{p}{PA}\rightsquigarrow{}_{p}{B}$ is simply the one obtained by restricting ${PA}\rightsquigarrow B$, since $FA\rightsquigarrow B$ restricts to $F{}_{p}{A}\rightsquigarrow{}_{p}{B}$. ∎
By construction, the sampling morphisms in $\mathsf{CPU}_{\min/\max}$ are simply given by the inclusions $x\mapsto[x]$, while the delta morphisms correspond to the multiplication maps
$[x_{1}\odot\dots\odot x_{n}]\longmapsto x_{1}\cdots x_{n}.$
In particular, although the multiplication map is usually not bounded as a map $A\odot A\rightsquigarrow A$ with respect to either the minimal or maximal tensor norm, we now have an equivalent formulation given by a bounded ***-homomorphism $\delta_{A}^{\mathrm{op}}\colon PA\rightsquigarrow A$.
The laxator $\nabla_{A,B}$ defined in (3.2) corresponds to the following map on simple tensors:
$[(x_{1}\odot y_{1})\odot\dots\odot(x_{n}\odot y_{n})]\longmapsto[x_{1}\odot\dots\odot x_{n}]\odot[y_{1}\odot\dots\odot y_{n}],$
where $x_{j}\in A$ and $y_{j}\in B$.
In the classical setting, observational representability was introduced in [16, Appendix A.4] based on the notion of observational monad from [34]. The idea behind this concept is that iterated sampling can distinguish any two distinct probability measures. Using 2.39, we can consider the morphisms $\mathsf{samp}^{(n)}_{A}\colon PA\to A^{\otimes n}$ for any $n\geq 1$.
Definition 3.37.
A picture ${\mathsf{D}}$ is observationally representable if it is representable and for every object $A$, the morphisms $(\mathsf{samp}_{A}^{(n)})_{n\geq 1}$ are jointly monic in $\operatorname{Gen}({\mathsf{D}})$: for any generalized morphisms $\phi,\psi\colon B\to PA$,
$\mathsf{samp}_{A}^{(n)}\phi=\mathsf{samp}_{A}^{(n)}\psi\quad\forall\,n\qquad\implies\qquad\phi=\psi.$
Remark 3.38.
If we assume that ${\mathsf{D}}$ has countable Kolmogorov products, then the family $\mathsf{samp}_{A}^{(n)}$ glues together to a single generalized morphism $\mathsf{samp}_{A}^{(\mathbb{N})}\colon PA\to A_{\mathbb{N}}$. The requirement of observational representability is then the same as saying that $\mathsf{samp}_{A}^{(\mathbb{N})}$ is monic in $\operatorname{Gen}({\mathsf{D}})$, since
$\mathsf{samp}_{A}^{(\mathbb{N})}\phi=\mathsf{samp}_{A}^{(\mathbb{N})}\psi\qquad\iff\qquad\mathsf{samp}_{A}^{(n)}\phi=\mathsf{samp}_{A}^{(n)}\psi\quad\forall\,n.$
Indeed the proof of this equivalence is an immediate application of the universal property of Kolmogorov products.
Proposition 3.39.
The pictures $\mathsf{CPU}_{\min/\max}$ are observationally representable.
Proof.
In view of Remark 3.38 and Proposition 3.27, we want to show that $\mathsf{samp}_{A}^{(\mathbb{N})}$ is monic for every $A$. This translates into proving that the map $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}\colon{}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{PA}$ is epic with respect to generalized morphisms.
We recall that ${}_{p}{PA}$ is a quotient of the free ***-algebra $F{}_{p}{A}$, as explained in the proof of Proposition 3.36. Therefore, the map $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ sends by construction an element $x_{1}\odot\dots\odot x_{n}\in{}_{p}{A^{\odot n}}\subseteq{}_{p}{A_{\mathbb{N}}}$ to the equivalence class of $x_{1}\odot\dots\odot x_{n}$ in ${}_{p}{PA}$. We conclude that $\mathsf{samp}_{A}^{(\mathbb{N}),\mathrm{op}}$ is surjective because every element in ${}_{p}{PA}$ is an equivalence class of finite sums of elements of the form $x_{1}\odot\dots\odot x_{n}$ for some $n$. ∎
Remark 3.40.
In $\mathsf{CPU}_{\min/\max}$, the iterated sampling morphism $\mathsf{samp}^{(\mathbb{N})}_{A}\colon PA\to A_{\mathbb{N}}$ is not self-adjoint in general. First, note that for self-adjoint elements $x,y\in{A}$, the element $x\odot y$ is self-adjoint in $A_{\mathbb{N}}$ while $[x\odot y]^{*}=[y\odot x]$ in $PA$. Therefore, $\mathsf{samp}^{(2)}$ is self-adjoint if and only if $PA$ is commutative: Indeed, the set of linear combinations of elements $x\odot y$ as above generates $PA$. If $A=0$, $\mathbb{C}$ or $\mathbb{C}^{2}$, then $PA$ is commutative: the first two are trivial because $P0=0$ and $P\mathbb{C}=\mathbb{C}$, while the third is discussed in [32, Section 5].
In the same section, Kirchberg and Wassermann prove that $P(\mathbb{C}^{3})$ and $P(M_{2})$ are not exact. This means in particular that they are not commutative. [^14] The exactness of a commutative C*-algebra is a standard fact. See e.g. [5] for more details (every commutative C*-algebra is nuclear, and every nuclear C*-algebra is exact). Therefore in these cases, the generalized morphisms $\mathsf{samp}^{(2)}$ and $\mathsf{samp}^{(\mathbb{N})}$ are not self-adjoint.
Proposition 3.41.
Let ${\mathsf{D}}$ be a representable picture. Then the following are equivalent:
  1. (i)
    ${\mathsf{D}}$ is observationally representable;
  2. (ii)
    For all objects $A$ and $B$, the morphisms $(samp^{(n)}_{A}\otimes\mathrm{id}_{B})_{n\in\mathbb{N}}$ are jointly monic in $\operatorname{Gen}({\mathsf{D}})$.
  3. (iii)
    For all objects $A$, the morphisms $(\mathsf{samp}^{(n)}_{A})_{n\in\mathbb{N}}$ are jointly monic in $\operatorname{Gen}({\mathsf{D}})$ modulo left a.s. equality: given any suitably composable generalized morphisms $\phi,\psi$ and $\omega$,
[…]
  4. (iv)
    For all objects $A$, the morphisms $(\mathsf{samp}^{(n)}_{A})_{n}$ are jointly monic in $\operatorname{Gen}({\mathsf{D}})$ modulo symmetric a.s. equality: given any suitably composable generalized morphisms $\phi,\psi,\omega$,
    $\mathsf{samp}_{A}^{(n)}\mathchoice{\,}{\,}{}{}\phi\cong_{\omega}\mathsf{samp}_{A}^{(n)}\mathchoice{\,}{\,}{}{}\psi\quad\forall\,n\qquad\implies\qquad\phi\cong_{\omega}\psi.$
By symmetry, we could equivalently formulate condition (iii) with respect to right a.s. equality.
Sketch of proof.
The implication (i)$\implies$(ii) is shown exactly as in the classical case [16, Proposition A.4.2]. Indeed, a direct computation shows that the coherence morphisms $\nabla_{A,B}$ are split monomorphisms exactly as expressed in [16, Lemma A.4.1]:
[…]
The other implications are simple checks. ∎
3.5 Classical representability
In this subsection, we consider a variation on the theme of representability which will turn out to provide a universal property for the state space of a C*-pair. This new representability is only required against classical objects, hence the name. The associated distribution objects, called state space objects, are obtained by the C*-algebra of complex-valued continuous functions on the state space. This fresh perspective on the state space is particularly fruitful. For example, the universal property implies that the C*-algebra of continuous functions on the state space is the abelianization of the universal C*-algebra, i.e. of the distribution object (Corollary 3.50).
Definition 3.42.
Let ${\mathsf{D}}$ be a picture.
  1. (i)
    Given an object $A$, a state space object is a classical object $SA$, together with an evaluation morphism $\mathsf{ev}_{A}\colon SA\to A$, such that the induced map
    $\mathsf{ev}_{A}\circ\_\colon{\mathsf{D}}_{\mathrm{det}}(B,SA)\to{\mathsf{D}}(B,A)$ (3.3)
    is bijective for every classical object $B\in{\mathsf{D}}$.
  2. (ii)
    We say that ${\mathsf{D}}$ is classically representable if every $A$ admits a state space object $SA$.
Notation 3.43.
We denote by $(\_)^{\natural}$ the inverse of (3.3). In particular, $\phi=\mathsf{ev}\mathchoice{\,}{\,}{}{}\phi^{\natural}$ for every morphism $\phi$ with classical domain, and every classical object $B$ will have a right inverse of $\mathsf{ev}_{B}$ given by $\mathrm{id}^{\natural}_{B}$.
Also, following Definition 2.37 (and Remark 3.6), we use the shorthands
$C\ell({\mathsf{D}})\coloneqq C\ell(\operatorname{Gen}({\mathsf{D}}))\cap{\mathsf{D}}\qquad\text{and}\qquad C\ell({\mathsf{D}})_{\mathrm{det}}\coloneqq C\ell({\mathsf{D}})\cap{\mathsf{D}}_{\mathrm{det}}.$
Classical representability has an analogue of Lemma 3.32.
Lemma 3.44.
Let ${\mathsf{D}}$ be a classically representable picture. Then $A\mapsto SA$ extends to a functor ${\mathsf{D}}\to C\ell({\mathsf{D}})_{\mathrm{det}}$ which is right adjoint to the inclusion $C\ell({\mathsf{D}})_{\mathrm{det}}\hookrightarrow{\mathsf{D}}$ with:
  1. (i)
    Unit given by the $\mathrm{id}^{\natural}$;
  2. (ii)
    Counit given by the evaluation morphisms $\mathsf{ev}$.
Remark 3.45.
The adjunction
[…]
induces a monad on $C\ell({\mathsf{D}})_{\mathrm{det}}$, which we also denote by $S$ by abuse of notation. In particular, for every $A$, the classical object $SA$ becomes an Eilenberg-Moore $S$-algebra with structure morphism $S(\mathsf{ev}_{A})$. This result is immediate from the monad construction. It is interesting to note that this $S$-algebra is not necessarily free, as we will see in Remark 3.48 below for $\mathsf{CPU}_{\min/\max}$.
Remark 3.46.
Let ${\mathsf{D}}$ be a classically representable picture having countable Kolmogorov products. For any object $A$, the state space object $S(A_{\mathbb{N}})$ is the limit of the diagram $\langle S(A_{F}),S(\pi_{F,F^{\prime}})\rangle$ indexed by finite $F\subseteq\mathbb{N}$ ordered under inclusion. Indeed, $S$ is a right adjoint by Lemma 3.44, and therefore preserves limits. This is a version of [42, Theorem 4.1] formulated in terms of pictures.
Proposition 3.47.
The pictures $\mathsf{CPU}_{\min/\max}$ are classically representable.
Proof.
For any C*-pair, a state is by definition a completely positive unital map ${}_{p}{A}\rightsquigarrow\mathbb{C}$, or equivalently $A\rightsquigarrow\mathbb{C}$. We write $\Sigma A$ for this set of states and equip it with the weak-* topology as usual, which is the weakest topology making the evaluation maps $\mathsf{ev}_{x}\colon\phi^{\mathrm{op}}\mapsto\phi^{\mathrm{op}}(x)$ continuous for every $x\in A$. It is well-known that $\Sigma{A}$ is a compact Hausdorff space. We now take
[…]
This is a commutative C*-algebra, which contains $\mathsf{ev}_{x}$ for every $x\in A$ by construction. We then define ${}_{p}{SA}$ to be the ***-subalgebra of $SA$ generated by $\mathsf{ev}_{x}$ for $x\in{}_{p}{A}$. It is dense since by the Hahn–Banach theorem, evaluation maps separate the points of $\Sigma{A}$, so that the density follows by the Stone–Weierstrass theorem. Next, we write
$\mathsf{ev}_{A}^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{SA}$
for the map sending each $x$ to its evaluation $\mathsf{ev}_{x}$. This map is clearly positive unital, hence completely positive by Proposition A.13. It is also injective: Indeed, whenever $\mathsf{ev}_{x}=\mathsf{ev}_{y}$, then $\omega^{\mathrm{op}}(x-y)=0$ for all states $\omega^{\mathrm{op}}$, so $x=y$ by Corollary A.19.
In order to show that ${SA}$ is the state space object, let us consider a classical C*-pair $(B,{}_{p}{B})$. Since the C*-algebra $B$ is commutative, by Gelfand duality it is ***-isomorphic to $C(X)$, the C*-algebra of complex-valued continuous map on a compact Hausdorff space $X$. In order to prove the desired bijection, we now consider for any morphism $\phi\colon B\to A$ the map
$\begin{array}[]{rcl}X&\to&\Sigma A\\ p&\mapsto&(x\mapsto\phi^{\mathrm{op}}(x)(p)),\end{array}$
which is continuous because post-composing with the evaluation maps $\mathsf{ev}_{x}$ gives continuous maps. By Gelfand duality, we therefore obtain a ***-homomorphism $\phi^{\natural,\mathrm{op}}\colon{}_{p}{SA}\rightsquigarrow{}_{p}{B}$ sending $\mathsf{ev}_{x}\in{}_{p}{SA}$ to $\phi^{\mathrm{op}}(x)$ by definition. This satisfies $\mathsf{ev}_{A}\mathchoice{\,}{\,}{}{}\phi^{\natural}=\phi$ by construction. Its uniqueness is immediate since this equation fixes $\phi^{\natural,\mathrm{op}}$ on the generators $\mathsf{ev}_{x}$ and $\phi^{\natural}$ is required to be a ***-homomorphism. ∎
Remark 3.48.
Let us work out what the $S$-algebra structure on state space objects $SA$ for the pictures $\mathsf{CPU}_{\min/\max}$ amounts to. As we will show, this is exactly the structure which equips the state spaces $\Sigma A$ with the structure of compact convex sets (cf. [22], and the explicit formulations [42, Theorems 2.18 and 2.19]). This is especially important in the foundations of quantum theory, where the convex structure is often considered fundamental due to its simple operational interpretation: a convex combination of states corresponds to a random ‘‘mixing’’ of the states that appear in it. [^15] To see that this was already recognized by von Neumann, note the importance he attributes to discussion of mixtures in his eminent treatise [47]. So although our formal categorical framework does not have any built-in notion of a convex structure, this comes out automatically from classical representability.
To begin, note that under Gelfand duality, there is a functor $C\ell(\mathsf{CPU}_{\min/\max})_{\mathrm{det}}\to\mathsf{CHaus}$ given by completion. In particular, if we restrict the objects to C*-algebras for simplicity, then this functor is an equivalence. On this subcategory, our monad $S$ corresponds to the Radon monad $P:\mathsf{CHaus}\to\mathsf{CHaus}$, which assigns to every compact Hausdorff space $X$ the space of Radon probability measures $PX$ [22]; this is straightforward to see from the proof of Proposition 3.47 by virtue of the natural isomorphism $\Sigma C(X)\cong PX$ given by the Riesz representation theorem. By a result of Świrszcz [44], the $P$-algebras are exactly the compact convex sets in locally convex spaces, where the structure morphism assigns to every Radon probability measure on such a set its barycenter. In particular, applying the structure morphism to finitely supported probability measures amounts to the formation of convex combinations. So in brief, the $P$-algebras are the compact convex sets. Its free algebras are the probability simplices, which are the closed convex hulls of the Dirac deltas on a compact Hausdorff space $X$ [^16] The precise notion of simplex here is that of Bauer simplex. See [1, Corollary II.4.2] for the theorem which identifies Bauer simplices with the spaces of Radon probability measures on compact Hausdorff spaces.
Now that we have clarified the relation between the Radon monad $P$ and our $S$, we can identify the $S$-algebra structure on $SA$ for every object $A$ in $\mathsf{CPU}_{\min/\max}$. An inspection of the proof of Proposition 3.47 shows that the structure morphism $S(\mathsf{ev}_{A})$ is given by
$\begin{array}[]{rcl}\Sigma C(\Sigma A)&\rightsquigarrow&\Sigma A\\ \mu&\mapsto&(a\mapsto\mu(\mathsf{ev}_{a})).\end{array}$
Under the isomorphism $\Sigma C(\Sigma A)\cong P(\Sigma A)$, this map becomes
$\mu\mapsto\mathopen{}\mathclose{{}\left(a\mapsto\int_{\Sigma A}\mathsf{ev}_{a}(s)\,\mu(\mathrm{d}s)}\right),$
which is exactly the barycenter formation on $\Sigma A$ considered as a compact convex set in the dual space $A^{*}$ equipped with the weak-* topology.
Finally, note that the state space $\Sigma A$ is generally not a simplex, and therefore not a free $P$-algebra. For example, $\Sigma M_{2}$ is famously a three-dimensional Euclidean ball, the Bloch ball [27]. More generally, the state space is a simplex if and only if $A$ is commutative [1, Corollary II.4.2]. Therefore any non-commutativity manifests itself in the non-uniqueness of the decomposition of a state into extremal states, a central theme in the foundations of quantum theory [37, Section 2.4.2].
Proposition 3.49 (Classicalization of the distribution object).
Assume that ${\mathsf{D}}$ is a representable and classically representable picture. Then
$\mathsf{ev}_{A}^{\sharp}\colon SA\to PA$
induces a bijection
$\mathsf{ev}^{\sharp}_{A}\circ\_\colon{\mathsf{D}}_{\mathrm{det}}(B,SA)\to{\mathsf{D}}_{\mathrm{det}}(B,PA)$
for every classical object $B$.
Proof.
Since $\mathsf{samp}_{A}\mathchoice{\,}{\,}{}{}\mathsf{ev}_{A}^{\sharp}=\mathsf{ev}_{A}$, we can consider the commutative diagram
[…]
which immediately proves the statement because the two arrows going to the right are bijections. ∎
We recall that the abelianization of a C*-algebra is defined as the quotient by the closed two-sided ideal generated by the commutators [3, Definition 2.8].
Corollary 3.50.
For every object $A$ in $\mathsf{CPU}_{\min/\max}$, the state space object ${SA}$ is the abelianization of the distribution object ${PA}$.
Proof.
Proposition 3.49 is exactly the universal property of the abelianization of ${PA}$ restricted to ***-homomorphisms, so the statement follows. ∎
Corollary 3.51.
For any objects $A$ and $B$ in $\mathsf{CPU}_{\min/\max}$, the morphism $\mathsf{ev}_{A}^{\sharp}\otimes\mathrm{id}_{B}$ is monic in $\operatorname{Gen}(\mathsf{CPU}_{\min/\max})$.
Proof.
By Corollary 3.50, we know that $\mathsf{ev}^{\sharp,\mathrm{op}}_{A}$ is surjective. A direct check shows that $\mathsf{ev}^{\sharp,\mathrm{op}}_{A}\otimes\mathrm{id}_{{B}}$ is also surjective. The statement follows since surjective maps give monomorphisms in the opposite category. ∎
Proposition 3.52.
Let ${\mathsf{D}}$ be a representable and classically representable picture. Then $(\mathsf{ev}_{A}^{\sharp})_{A\in{\mathsf{D}}}$ is a natural transformation $S\to P$, where $S$ and $P$ are both considered with target in ${\mathsf{D}}_{\mathrm{det}}$.
Proof.
Let $\phi\colon A\to B$ be any morphism in ${\mathsf{D}}$. We need to show that
[…]
Since all these morphisms are deterministic, it is sufficient to show that the diagram commutes after post-composing with $\mathsf{samp}_{B}$. This is a direct computation:
$\mathsf{samp}_{B}\mathchoice{\,}{\,}{}{}\mathsf{ev}^{\sharp}_{B}\mathchoice{\,}{\,}{}{}S(\phi)=\mathsf{ev}_{B}\mathchoice{\,}{\,}{}{}S(\phi)=\phi\mathchoice{\,}{\,}{}{}\mathsf{ev}_{A}=\phi\mathchoice{\,}{\,}{}{}\mathsf{samp}_{A}\mathchoice{\,}{\,}{}{}\mathsf{ev}_{A}^{\sharp}=\mathsf{samp}_{B}\mathchoice{\,}{\,}{}{}P(\phi)\mathchoice{\,}{\,}{}{}\mathsf{ev}_{A}^{\sharp}.\qed$
Definition 3.53.
A picture ${\mathsf{D}}$ is observationally classically representable if it is classically representable and for all objects $A$ and $B$, the morphisms $(\mathsf{ev}_{A}^{(n)}\otimes\mathrm{id}_{B})_{n\geq 1}$ are jointly monic in $\operatorname{Gen}({\mathsf{D}})$: For all suitably composable generalized morphisms $\phi$ and $\psi$,
$(\mathsf{ev}_{A}^{(n)}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\phi=(\mathsf{ev}_{A}^{(n)}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\psi\quad\forall\,n\qquad\implies\qquad\phi=\psi.$
This definition is motivated by the fact that Items (iii) and (iv) of Proposition 3.41 now also hold with $\mathsf{ev}^{(n)}$ in place of $\mathsf{samp}^{(n)}$. It is not clear to us whether leaving out the extra tensor factor $B$ from the definition would still result in an analogue of Proposition 3.41.
Lemma 3.54.
Let ${\mathsf{D}}$ be an observationally representable picture, which is also classically representable. The following assertions are equivalent:
  • The picture ${\mathsf{D}}$ is observationally classically representable;
  • For all objects $A$ and $B$, the morphism $\mathsf{ev}^{\sharp}_{A}\otimes\mathrm{id}_{B}$ is monic in $\operatorname{Gen}({\mathsf{D}})$.
Proof.
Since $\mathsf{ev}^{(n)}_{A}\otimes\mathrm{id}_{B}=(\mathsf{samp}^{(n)}_{A}\mathchoice{\,}{\,}{}{}\mathsf{ev}_{A}^{\sharp})\otimes\mathrm{id}_{B}$, by Proposition 3.41(ii) we obtain that
$(\mathsf{ev}^{(n)}_{A}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\phi=(\mathsf{ev}^{(n)}_{A}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\psi$
for all $n$ if and only if $(\mathsf{ev}^{\sharp}_{A}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\phi=(\mathsf{ev}^{\sharp}_{A}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}\psi$. The statement follows. ∎
Proposition 3.55.
The pictures $\mathsf{CPU}_{\min/\max}$ are observationally classically representable.
Proof.
By Corollary 3.51, $\mathsf{ev}_{A}^{\sharp}\otimes\mathrm{id}_{B}$ is monic, so we can apply Lemma 3.54 to conclude the proof for $\mathsf{CPU}_{\min/\max}$. ∎
Concerning representability, we know that every $\mathsf{samp}_{A}$ is epic because it has a right inverse given by $\delta_{A}$. It is therefore natural to ask whether $\mathsf{ev}_{A}$ is also epic. This is indeed the case, provided we make the additional assumption of local state-separability. We first generalize the definition from [17, Definition 2.2.3(i)].
Definition 3.56.
A picture ${\mathsf{D}}$ is locally state-separable if the following implication holds for all generalized morphisms $\phi,\psi\colon W\otimes A\to B$:
[…]
Proposition 3.57.
The pictures $\mathsf{CPU}_{\min/\max}$ are locally state-separable.
The proof is immediate from the following result, with notation matching the one of (3.4).
Lemma 3.58.
Let $A$ and $W$ be C*-pairs. If $x\in{W}\otimes{A}$ satisfies $(\omega^{\mathrm{op}}\otimes\mathrm{id}_{A})(x)=0$ for every state ${\omega}{}\colon\mathbb{C}\to{W}$, then $x=0$.
Proof.
We consider $x\in W\otimes A$ and choose a decomposition $x=\sum_{i}w_{i}\otimes a_{i}$, where the $a_{i}$ are linearly independent. Then
$0=(\omega^{\mathrm{op}}\otimes\mathrm{id}_{A})(x)=\sum_{i}\omega^{\mathrm{op}}(w_{i})a_{i}$
implies that $\omega^{\mathrm{op}}(w_{i})=0$ for all $i$. We are therefore reduced to showing that whenever $\omega^{\mathrm{op}}(w_{i})=0$ for all states $\omega^{\mathrm{op}}$, then $w_{i}=0$. This is Corollary A.19. ∎
Remark 3.59.
In the classical case, local state-separability was ensured as a consequence of point-separability [16, Lemma 2.2.4]. Although the proof of this implication is still valid in our setting, this is not of much interest as point-separability is not satisfied for $\mathsf{CPU}_{\min/\max}$, because there are not sufficiently many deterministic states. For instance, for $n\geq 2$ there is no multiplicative linear map $M_{n}\rightsquigarrow\mathbb{C}$, because $M_{n}$ is a simple noncommutative algebra. Therefore, choosing two distinct morphisms with the same domain and target $M_{n}$ gives a counterexample to point-separability.
Lemma 3.60.
Let ${\mathsf{D}}$ be a classically representable locally state-separable picture. Then $\mathsf{ev}_{A}\otimes\mathrm{id}_{B}$ is epic in $\operatorname{Gen}(D)$ for all objects $A$ and $B$. Moreover, $S\colon{\mathsf{D}}\to C\ell({\mathsf{D}})_{\mathrm{det}}$ is faithful.
Proof.
Let $\phi,\psi\colon A\otimes B\to C$ be two morphisms such that $\phi\mathchoice{\,}{\,}{}{}(\mathsf{ev}_{A}\otimes\mathrm{id}_{B})=\psi\mathchoice{\,}{\,}{}{}(\mathsf{ev}_{A}\otimes\mathrm{id}_{B})$. By local state-separability, $\phi=\psi$ iff $\phi\mathchoice{\,}{\,}{}{}(\omega\otimes\mathrm{id}_{B})=\psi\mathchoice{\,}{\,}{}{}(\omega\otimes\mathrm{id}_{B})$ for every state $\omega\colon I\to A$. Since $I$ is classical, classical representability shows that $\omega$ factors through $\mathsf{ev}_{A}$. Finally, we have
[…]
for all $\omega$, so that $\phi=\psi$.
The last sentence is a general fact about adjoint functors: Faithfulness of the right adjoint is equivalent to the counit being epic. ∎
3.6 The quantum de Finetti theorem
In classical probability theory, the de Finetti theorem states that every permutation-invariant distribution of an infinite sequence of random variables is given by a mixture of i.i.d. distributions. The fact that this representation is unique amounts to a universal property characterizing distribution objects. In this way, the de Finetti theorem is closely related to representability.
In this subsection, we consider these questions in the involutive setting. The relevant flavor of representability is classical representability, where the role of distribution objects is now played by state space objects. One theoretical motivation for this is that we would like to factor through an iterated sampling map, and such an operation should be physically implementable, which is not satisfied for general representability (see Remark 3.40). Based on the quantum de Finetti theorem for states proven as Theorem C.2, we note that these properties apply in our pictures of interest $\mathsf{CPU}_{\min/\max}$.
Definition 3.61.
Let ${\mathsf{D}}$ be a picture with countable Kolmogorov products. A morphism $\phi\colon C\to A_{\mathbb{N}}\otimes B$ is exchangeable in the first factor if, for any finite permutation $\sigma\colon\mathbb{N}\to\mathbb{N}$, we have
[…]
If $B$ is omitted, or equivalently $B=I$, such $\phi$ is called exchangeable.
Here, we use the double wire to indicate an infinite Kolmogorov product [18], while $A_{\sigma}$ is as in 3.28.
Definition 3.62.
Let ${\mathsf{D}}$ be a picture with countable Kolmogorov products. For an object $A$, a de Finetti object is an object $QA$ together with a morphism $\ell_{A}\colon QA\to A$ such that:
  1. (i)
    $\ell_{A}$ is autocompatible;
  2. (ii)
    $\ell_{A}^{(\mathbb{N})}\colon QA\to A_{\mathbb{N}}$ is the universal morphism making $QA$ into the equalizer of all finite permutations
    $A_{\sigma}\colon A_{\mathbb{N}}\to A_{\mathbb{N}}$
    in ${\mathsf{D}}$, and this limit is preserved by every _$\_\otimes\mathrm{id}_{B}$ for every object $B$;
  3. (iii)
    $\ell_{A}^{(\mathbb{N})}\otimes\mathrm{id}_{B}$ is monic in $\operatorname{Gen}({\mathsf{D}})$.
In more detail, Property (ii) says the following: every morphism $\phi\colon C\to A_{\mathbb{N}}\otimes B$ exchangeable in the first factor admits a decomposition of the form
[…]
for a unique morphism $\mu\colon C\to QA\otimes B$.
Remark 3.63.
An alternative and categorically more elegant description of Property (ii) of Definition 3.62 is obtained by using a different diagram, given by all the finite tensor powers $A_{F}$ as in Definition 3.24, together with all those deterministic morphisms $A_{F}\to A_{F^{\prime}}$ that are induced from injections $F^{\prime}\hookrightarrow F$. In this way, one can make sense of the universal property of a de Finetti object without requiring the existence of countable Kolmogorov products. Such an approach is also more consistent with the limit considered by Staton and Summers [42, Lemma 4.2 and Theorem 4.3]. However, Property (ii) seems more natural from an intuitionistic point of view, as it fits better with the narrative of exchangeable morphisms.
Here is now our main categorical result.
Theorem 3.64.
Let ${\mathsf{D}}$ be a picture with countable Kolmogorov products. If $A$ has a de Finetti object, then $QA$ is a state space object with evaluation morphism $\ell_{A}$.
This is somewhat surprising, since we did not explicitly ask anything about the behavior of $QA$ with respect to deterministic morphisms.
Proof.
First, let us show that $QA$ is classical. Indeed, by Corollary 2.45, the autocompatibility of $\ell_{A}$ itself and the definition of Kolmogorov products, $\ell_{A}^{(\mathbb{N})}$ is autocompatible. Property (iii) of Definition 3.62 now reduces the classicality of $QA$ to an application of Lemma 2.46. Second, we aim to show that $\ell_{A}\colon QA\to A$ gives a bijection as expressed in (3.3). So let us fix any classical object $B$. Then the injectivity of
$\ell_{A}\circ\_\colon{\mathsf{D}}_{\mathrm{det}}(B,QA)\to{\mathsf{D}}(B,A)$
is quite clear: If $\ell_{A}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}\mathchoice{\,}{\,}{}{}\psi$, then also $\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\phi=\ell_{A}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\psi$ because $\phi$ and $\psi$ are deterministic, and so $\phi=\psi$ by monicness of $\ell_{A}^{(\mathbb{N})}$. To prove surjectivity, consider a morphism $\phi\colon B\to A$. The classicality of $B$ implies that $\phi$ is autocompatible (Remark 2.41(i)), and therefore $\phi^{(\mathbb{N})}\colon B\to A_{\mathbb{N}}$ belongs to ${\mathsf{D}}$ as mentioned in Remark 3.26. By the universal property, we have $\phi^{(\mathbb{N})}=\ell^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\mu$ for some $\mu\colon B\to QA$. By marginalization, we immediately get $\phi=\ell\mathchoice{\,}{\,}{}{}\mu$, and hence
[…]
Using this equality, together with the definition of Kolmogorov products and associativity of $\operatorname{copy}_{B}$, we obtain
[…]
Since $\ell^{(\mathbb{N})}\otimes\mathrm{id}_{A}$ is monic, the same is true for $\ell^{(\mathbb{N})}\otimes\ell^{(\mathbb{N})}$. Therefore, $\mu$ is deterministic, and surjectivity is ensured. ∎
In view of this result, de Finetti objects will be denoted by $(SA,\mathsf{ev})$. We also obtain a new notion of representability.
Definition 3.65.
Let ${\mathsf{D}}$ be a picture. Then ${\mathsf{D}}$ is de Finetti representable if it has countable Kolmogorov products and a de Finetti object for every object.
Due to Property (iii) of Definition 3.62, an immediate consequence of Theorem 3.64 is the following.
Corollary 3.66.
Any de Finetti representable picture is observationally classically representable.
It is worth noting that there is a strong connection between this result and the recent proof of the de Finetti theorem for Markov categories [18, Theorem 4.4]. However, we refrain from discussing this here, as it is beyond the scope of the present paper.
Lemma 3.67.
Let ${\mathsf{D}}$ be a de Finetti representable picture. For every pair of objects $A$ and $B$, we have a natural bijection
${\mathsf{D}}_{\mathrm{det}}(B,SA)\cong\mathopen{}\mathclose{{}\left\{\textrm{Autocompatible morphisms }B\to A}\right\}$
given by composition with $\mathsf{ev}_{A}$.
For $\mathsf{CPU}_{\min/\max}$ this result is expected: autocompatible morphisms translate to maps with commutative image (see Example 2.42), and therefore they factor through a deterministic epimorphism with classical codomain. However, de Finetti representability allows an elegant highway independent of such an explicit factorization.
Proof.
The reasoning is exactly the same as employed in the proof of Theorem 3.64. Indeed, the fact that $B$ was classical ensured that, given $\phi\colon B\to A$, we could construct an exchangeable $\phi^{(\mathbb{N})}$, and this is exactly the requirement of $\phi$ to be autocompatible. The only thing that may be unclear is why $\mathsf{ev}\mathchoice{\,}{\,}{}{}\phi$, with $\phi\colon B\to SA$ deterministic, is autocompatible. This follows from the fact that $\phi$ is deterministic and $SA$ is classical: $(\mathsf{ev}\mathchoice{\,}{\,}{}{}\phi)^{(\mathbb{N})}=\mathsf{ev}^{(\mathbb{N})}\mathchoice{\,}{\,}{}{}\phi$, and the latter is a composition of two morphisms in ${\mathsf{D}}$. ∎
We now focus our attention on our pictures of interest. Let us start with a technical lemma.
Lemma 3.68.
Let $\phi^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{A}^{\prime}$ be any surjective positive unital map. Then for any C*-pair $B$, composing with $\phi$ establishes a bijection between
  1. (i)
    the set of positive unital maps ${}_{p}{A}^{\prime}\rightsquigarrow{}_{p}{B}$, and
  2. (ii)
    the set of positive unital maps $\psi^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{B}$ such that $\mathopen{}\mathclose{{}\left\lVert\psi^{\mathrm{op}}(x)}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}(x)}\right\rVert$ for every self-adjoint element $x$.
Proof.
Given any positive unital map $\psi^{\prime\mathrm{op}}\colon{}_{p}{A^{\prime}}\rightsquigarrow{}_{p}{B}$, we clearly have
$\mathopen{}\mathclose{{}\left\lVert\psi^{\prime\mathrm{op}}\mathchoice{\,}{\,}{}{}\phi^{\mathrm{op}}(x)}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}(x)}\right\rVert$
for every $x\in{}_{p}{A}$, so that the map is well-defined. The injectivity is obvious because $\phi^{\mathrm{op}}$ is surjective.
For surjectivity, suppose that $\psi^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{B}$ satisfies the given inequality. Then for any self-adjoint $x\in{}_{p}{A}$, we can choose any self-adjoint [^17] Since $x$ has some preimage, it also has a self-adjoint preimage, namely the self-adjoint part of any preimage. $y\in{}_{p}{A^{\prime}}$ with $\phi^{\mathrm{op}}(y)=x$ and define $\psi^{\prime\mathrm{op}}(x)\coloneqq\psi^{\mathrm{op}}(y)$. This is well-defined since $\ker(\phi^{\mathrm{op}})\subseteq\ker(\psi^{\mathrm{op}})$ by the assumed inequality, and this also implies the linearity of $\psi^{\prime\mathrm{op}}$ on the self-adjoint part of $\operatorname{\operatorname{im}}(\phi^{\mathrm{op}})$. Furthermore, the assumed inequality ensures that $\mathopen{}\mathclose{{}\left\lVert\psi^{\prime\mathrm{op}}(x)}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}(x)}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x}\right\rVert$, where the second inequality is by Proposition A.9. It follows that $\psi^{\prime\mathrm{op}}$ extends to a bounded unital map on $\operatorname{\operatorname{im}}(\phi^{\mathrm{op}})$, which is equal to ${}_{p}{A^{\prime}}$ as noted above. This map has norm $\leq 1$ on self-adjoints, and is therefore positive by Corollary A.20. ∎
Theorem 3.69 (Quantum de Finetti, categorical version).
The pictures $\mathsf{CPU}_{\min/\max}$ are de Finetti representable. In particular, for any morphism $\phi\colon C\to A_{\mathbb{N}}\otimes B$ exchangeable in the first factor, there exists a unique morphism $\mu\colon C\to SA\otimes B$ such that
[…]
Proof.
We first construct a candidate de Finetti object $QA$ for these pictures and then prove that it indeed has the required properties. The fact that it is naturally isomorphic to the state space object $SA$, as indicated in the second part of the claim, is then clear by Theorem 3.64.
To start, we first define $QA$ as a C*-algebra before describing the dense subalgebra ${}_{p}{QA}$ and the morphism $\ell\colon A_{\mathbb{N}}\to QA$. This works similarly as for the construction of $PA$ in the proof of Proposition 3.36. We equip the Kolmogorov power $A_{\mathbb{N}}$ with the seminorm
[…]
Then by exchangeability, it is clear that for every finite permutation $\sigma$, the elements of the form $x-A_{\sigma}^{\mathrm{op}}(x)$ have norm zero. In particular, defining $QA$ as the completion of the quotient of $A_{\mathbb{N}}$ by the nullspace $\mathcal{N}_{QA}$ of this seminorm makes the projection map $A_{\mathbb{N}}\rightsquigarrow QA$ invariant under finite permutations. The fact that elements supported on finitely many tensor factors are dense in $A_{\mathbb{N}}$ and that the quotient map $A_{\mathbb{N}}\rightsquigarrow QA$ has norm $\leq 1$ shows that it is invariant even under all permutations $\mathbb{N}\to\mathbb{N}$.
Therefore we can define a multiplication on $QA$ by ‘‘disjoint tensoring’’ as follows. For $x,y\in A_{\mathbb{N}}$, we have $x\otimes y\in A_{\mathbb{N}+\mathbb{N}}$, and we can consider this as an element of $A_{\mathbb{N}}$ via any bijection $\mathbb{N}+\mathbb{N}\cong\mathbb{N}$, for which we also write $x\otimes y$ by abuse of notation. We then define the multiplication on $QA$ as
$[x]\cdot[y]\coloneqq[x\otimes y],$
and the permutation invariance implies that the result is independent of the choice of bijection. For well-definedness, we first note that by Bauer’s maximum principle [4, Theorem 1.4.1] and the analytic quantum de Finetti theorem for states (Theorem C.2), we can equivalently write
[…]
Therefore if $\mathopen{}\mathclose{{}\left\lVert x}\right\rVert_{QA}=0$ or $\mathopen{}\mathclose{{}\left\lVert y}\right\rVert_{QA}=0$, then also $\mathopen{}\mathclose{{}\left\lVert x\otimes y}\right\rVert_{QA}=0$ for any $y\in A_{\mathbb{N}}$, and hence the multiplication is well-defined on equivalence classes. The star operation on $A_{\mathbb{N}}$ similarly descends to an antilinear map $QA\rightsquigarrow QA$, and it is now straightforward to see that $QA$ is a ***-algebra.
The submultiplicativity of the norm and the C*-identity on $QA$ also follow easily from (3.6), and therefore $QA$ is a C*-algebra. We define the dense ***-subalgebra as the image of ${}_{p}{A_{\mathbb{N}}}$ under the quotient map, and take $\ell_{A}^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ to be the composition of the inclusion ${}_{p}{A}\rightsquigarrow{}_{p}{A_{\mathbb{N}}}$ in the first factor and the quotient map ${}_{p}{A_{\mathbb{N}}}\rightsquigarrow{}_{p}{QA}$ itself. Then $\ell_{A}^{\mathrm{op}}$ is unital and contractive, and therefore positive by Proposition A.9. Moreover since $QA$ is commutative, complete positivity follows by Proposition A.13, and also $\ell_{A}$ is trivially autocompatible. Thus Property (i) of Definition 3.62 is satisfied. The monicness of $\ell_{A}\otimes\mathrm{id}_{B}$ for every $B$ required by Property (iii) is clear, since $\ell_{A}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}$ surjectively onto ${}_{p}{QA}$, and hence also $\ell_{A}\otimes\mathrm{id}_{B}^{\mathrm{op}}$ maps ${}_{p}{A_{\mathbb{N}}}\odot{}_{p}{B}$ surjectively onto ${}_{p}{QA}\odot{}_{p}{B}$.
Finally for Property (ii), let us consider the case of positive unital maps first. Then by Lemma 3.68 it is enough to show that for every positive unital $\rho\colon C\to A_{\mathbb{N}}\otimes B$ exchangeable in the first factor, we have
$\mathopen{}\mathclose{{}\left\lVert\rho^{\mathrm{op}}(x)}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert(\ell_{A}\otimes\mathrm{id}_{B})^{\mathrm{op}}(x)}\right\rVert$
for all self-adjoint $x\in A_{\mathbb{N}}\otimes B$. Since $\rho^{\mathrm{op}}(x)$ is also self-adjoint, we have
[…]
by Corollary A.17. But since $\rho\mathchoice{\,}{\,}{}{}\eta\colon\mathbb{C}\to A_{\mathbb{N}}\otimes B$ is a state exchangeable in the first factor, it only remains to be shown that
$\sup_{\phi}\,\mathopen{}\mathclose{{}\left\lvert\phi^{\mathrm{op}}(x)}\right\rvert\leq\mathopen{}\mathclose{{}\left\lVert(\ell_{A}\otimes\mathrm{id}_{B})^{\mathrm{op}}(x)}\right\rVert$ (3.7)
where $\phi\colon\mathbb{C}\to A_{\mathbb{N}}\otimes B$ ranges over states exchangeable in the first factor. The case $B=\mathbb{C}$ is immediate by definition of _$\mathopen{}\mathclose{{}\left\lVert\_}\right\rVert_{QA}$. In general, by another application of Bauer’s maximum principle and the quantum de Finetti theorem for states (Theorem C.2),
[…]
The case $B=\mathbb{C}$ shows that any power state $\psi^{(\mathbb{N})}$ factors through $\ell^{(\mathbb{N})}$: let $\mu$ be the state on $QA$ such that $\psi^{(\mathbb{N})}=\ell^{(\mathbb{N})}\mu$. In particular, $(\psi^{(\mathbb{N})},\omega)=(\ell^{(\mathbb{N})}\otimes\mathrm{id}_{B})\mathchoice{\,}{\,}{}{}(\mu,\omega)$. An application of Corollary A.17 is then enough to conclude that (3.7) holds.
Property (ii) for $\mathsf{CPU}_{\min/\max}$ now follows from the case of positive unital maps, given that we can simply add an additional tensor factor to both $B$ and $C$. ∎
The argument used in the proof of Theorem 3.69 shows that the factorization is in fact more general, i.e. it holds for any $n$-positive [^18] See e.g. [40, p. 26] for a definition. unital map which is exchangeable in the first factor, for any choice of positive integer $n$. In particular, by Proposition A.13 we obtain the following immediate consequence.
Corollary 3.70.
An exchangeable positive unital map $\phi^{\mathrm{op}}\colon{}_{p}{A}_{\mathbb{N}}\rightsquigarrow{}_{p}{B}$, where the tensor is with respect to either the maximal or the minimal tensor norm, is completely positive.
Combining Lemmas 3.67 and 3.69, we also get the following.
Corollary 3.71.
Given any two C*-pairs $A$ and $B$, we have natural bijections
$\mathsf{CPU}_{\min/\max,\mathrm{det}}(B,SA)\cong\mathopen{}\mathclose{{}\left\{\textrm{Autocompatible morphisms }B\to A}\right\}.$
Appendix A From C*-algebras to C*-pairs
Here we recall some basic definitions and results on C*-algebras and indicate how these generalize to C*-pairs. A first step in this direction is to relate the latter to pre-C*-algebras, i.e. C*-algebras that are not necessarily complete. This is Proposition A.3. We then ensure that the relevant standard theory from the C*-setting still works if our algebras are not complete, provided that the definitions are suitably adapted. Of course, not all results on Banach spaces extend to normed vector spaces: For example, the open mapping theorem and the closed graph theorem are two well-known examples of statements that fail to hold without assuming completeness. For our purposes, however, virtually everything works without completeness. The only exception is a detail on boundedness, namely Example 3.13.
C*-pairs and pre-C*-algebras
For us, all algebras are assumed unital without further mention, and all ***-homomorphisms are assumed to preserve the unit. To distinguish between the algebraic tensor product and the tensor product of C*-algebras, we use $\odot$ and $\otimes$, respectively (see 2.30).
Definition A.1.
A pre-C*-algebra ${}_{p}{A}$ is a normed ***-algebra, i.e. a ***-algebra [^19] For a brief recap of what a ***-algebra is, see Example 2.25. with a submultiplicative norm, satisfying the C*-identity
$\mathopen{}\mathclose{{}\left\lVert x^{*}x}\right\rVert=\mathopen{}\mathclose{{}\left\lVert x}\right\rVert^{2}$
for every $x\in{}_{p}{A}$.
The decoration ${}_{p}{\,}$ is used to emphasize that we are dealing with pre-C*-algebras instead of C*-algebras. If the norm on ${}_{p}{A}$ is complete, we omit the decoration and say that $A$ is a C*-algebra.
Lemma A.2.
The completion $\operatorname{cl}({{}_{p}{A}})$ of a pre-C*-algebra ${}_{p}{A}$ is a C*-algebra. Conversely, any ***-subalgebra of a C*-algebra is a pre-C*-algebra with respect to the induced norm.
Proof.
Since the addition, the product and the involution respect limits, the completion becomes a C*-algebra. The second part is immediate. ∎
Proposition A.3.
The category $\operatorname{C*-}{\mathsf{pr}}$ (Definition 2.27) is equivalent to the category with
  • Pre-C*-algebras as objects,
  • Formal opposites of unital linear maps as morphisms,
in such a way that the associated natural isomorphisms have isometric ***-isomorphisms as components.
Proof.
By Lemma A.2, every dense ***-subalgebra of a C*-algebra is a pre-C*-algebra. Then the functor $(A,{}_{p}{A})\mapsto{}_{p}{A}$ is clearly fully faithful and essentially surjective, so the two categories are equivalent. Its inverse is given by sending ${}_{p}{A}$ to the pair $(\operatorname{cl}({{}_{p}{A}}),{}_{p}{A})$. Therefore ${}_{p}{A}\mapsto(\operatorname{cl}({{}_{p}{A}}),{}_{p}{A})\mapsto{}_{p}{A}$ is precisely the identity functor, while $(A,{}_{p}{A})\mapsto{}_{p}{A}\mapsto(\operatorname{cl}({{}_{p}{A}}),{}_{p}{A})$ is naturally isomorphic to the identity functor on $\operatorname{C*-}{\mathsf{pr}}$ via the isomorphisms $(A,{}_{p}{A})\cong(\operatorname{cl}({{}_{p}{A}}),{}_{p}{A})$ implemented by the identities ${}_{p}{A}\rightsquigarrow{}_{p}{A}$. ∎
Although this was not needed in the proof, it is worth noting that for every C*-pair $(A,{}_{p}{A})$, the C*-algebra $A$ is canonically ***-isomorphic to the completion $\operatorname{cl}({{}_{p}{A}})$. For this reason, we will now simply write $A$ for the completion of ${}_{p}{A}$.
Lemma A.4.
Let $(A,{}_{p}{A})$ and $(B,{}_{p}{B})$ be C*-pairs and $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ a bounded linear map. Then $\phi^{\mathrm{op}}$ has a unique bounded extension $\operatorname{cl}({\phi^{\mathrm{op}}})\colon{B}\rightsquigarrow{A}$ between C*-algebras, and $\mathopen{}\mathclose{{}\left\lVert\operatorname{cl}({\phi^{\mathrm{op}}})}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}}\right\rVert$.
Proof.
For $x=\lim_{n}x_{n}$ with $x\in{}_{p}{B}$, we define $\operatorname{cl}({\phi^{\mathrm{op}}})(x)\coloneqq\lim_{n}\phi^{\mathrm{op}}(x_{n})$. Since $\phi^{\mathrm{op}}$ is bounded, this is well-defined. Uniqueness is immediate since for every limit a bounded map satisfies $\phi^{\mathrm{op}}(\lim_{n}x_{n})=\lim_{n}\phi^{\mathrm{op}}(x_{n})$. The equation $\mathopen{}\mathclose{{}\left\lVert\operatorname{cl}({\phi^{\mathrm{op}}})}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}}\right\rVert$ holds by continuity of the norm. ∎
Proposition A.5.
Let $(A,{}_{p}{A})$ and $(B,{}_{p}{B})$ be C*-pairs. Then ${}_{p}{A}\odot{}_{p}{B}$ is dense in $A\otimes B$, the C*-algebra obtained by completing $A\odot B$ with respect to any C*-norm, i.e. a submultiplicative norm satisfying the C*-identity.
Proof.
Clearly $A\odot B$ is dense in $A\otimes B$ by definition of the latter, so it is enough to show that ${}_{p}{A}\odot{}_{p}{B}$ is dense in $A\odot B$. For a simple tensor $x\odot y\in A\odot B$, we have that $x=\lim_{n}x_{n}$ and $y=\lim_{m}y_{m}$ are limits of sequences $(x_{n})$ in ${}_{p}{A}$ and $(y_{m})$ in ${}_{p}{B}$, respectively. Then a standard argument shows that $x\odot y=\lim_{n}{x_{n}\odot y_{n}}$ by submultiplicativity of the norm. Since any element of $A\odot B$ is a sum of simple tensors, and limits commute with addition, we conclude that ${}_{p}{A}\odot{}_{p}{B}$ is dense in $A\odot B$, as claimed. ∎
Proposition A.6 (Gelfand duality).
Let us consider
  • ${\mathsf{CHaus}}$, the category of compact Hausdorff spaces with continuous maps, and
  • $\operatorname{CC*-}{\mathsf{alg}}_{\mathrm{det}}$, the category of commutative C*-algebras with (bounded) ***-homomorphisms.
Then there is a contravariant equivalence
${\mathsf{CHaus}}\xrightarrow{\,\cong\,}\operatorname{CC*-}{\mathsf{alg}}_{\mathrm{det}}$
sending each compact Hausdorff space $X$ to the commutative C*-algebra $C(X)$ of complex-valued continuous maps.
For more details on this result, the reader may refer to [31, Theorem 1.1.1].
Positivity in pre-C*-algebras
Before dealing with positive maps, let us recall what a positive element is. Positivity is defined with respect to the containing C*-algebra.
Definition A.7.
For a C*-pair $(A,{}_{p}{A})$, we say that $x\in{}_{p}{A}$ is
  1. (i)
    self-adjoint if $x=x^{*}$;
  2. (ii)
    positive, and we write $x\geq 0$, if $x=y^{*}y$ for some $y\in A$.
Each element is a linear combination of two self-adjoints:
$x=\frac{x+x^{*}}{2}+\mathsf{i}\,\frac{x-x^{*}}{2\mathsf{i}},$
where $\mathsf{i}$ is the imaginary unit. Moreover, positivity induces a partial order, with $x\geq y$ short for $x-y\geq 0$.
Definition A.8.
Let ${}_{p}{A}$ and ${}_{p}{B}$ be pre-C*-algebras. Then a unital linear map $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ is
  • self-adjoint if $\phi^{\mathrm{op}}(x^{*})=\phi^{\mathrm{op}}(x)^{*}$;
  • positive if $\phi^{\mathrm{op}}(x)\geq 0$ for all $x\geq 0$;
  • completely positive if $\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}$ is positive [^20] One may wonder if we should emphasize which C*-norm is used; this is luckily not the case because $M_{n}$ is a nuclear C*-algebra (see, for example, [48, Proposition T.5.20]). for all $n$ ($M_{n}$ is the C*-algebra of matrices $n\times n$ with complex entries).
A positive unital functional $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow\mathbb{C}$ is called state.
Proposition A.9.
The positive unital maps ${}_{p}{B}\rightsquigarrow{}_{p}{A}$ are precisely the linear unital maps of norm $\leq 1$, and they are self-adjoint.
We refer to [40] for the proofs and more details, in particular Exercise 2.1, Corollary 2.9 and Proposition 2.11 in there. The fact that boundedness is automatic enters crucially in the following proof.
Proposition A.10.
A unital map $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ is positive or completely positive if and only if $\operatorname{cl}({\phi^{\mathrm{op}}})\colon B\rightsquigarrow A$ is.
Proof.
The ‘‘if’’ direction is clear. For the ‘‘only if’’, let us consider a positive element $x=y^{*}y$ of ${{B}}$. Since $y\in{{B}}$, there is a convergent sequence $y_{n}\to y$ with $y_{n}\in{}_{p}{B}$ for all $n$. In particular, this means that $y_{n}^{*}y_{n}\to x$. By definition of $\operatorname{cl}({\phi^{\mathrm{op}}})$, we have $\operatorname{cl}({\phi^{\mathrm{op}}})(x)=\lim_{n}{\phi^{\mathrm{op}}}(y_{n}^{*}y_{n})$. Since the set of positive elements is closed [33, Theorem 1.4.8], we conclude that $\operatorname{cl}({\phi^{\mathrm{op}}})(x)$ is still positive.
In order to deal with complete positivity, let us first show that $\operatorname{cl}({\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}})=\operatorname{cl}({\phi^{\mathrm{op}}})\odot\mathrm{id}_{M_{n}}$. Indeed, $M_{n}$ is nuclear and the algebraic tensor products with $M_{n}$ are already complete [48, Proposition T.5.20]. This means that $\operatorname{cl}({\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}})\colon{B}\odot M_{n}\rightsquigarrow{A}\odot M_{n}$. In particular, both $\operatorname{cl}({\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}})$ and $\operatorname{cl}({\phi^{\mathrm{op}}})\odot\mathrm{id}_{M_{n}}$ extend $\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}$, so they must coincide, according to Lemma A.4. We now can use the first part of this proof also for complete positivity: indeed, $\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}$ is positive if and only if $\operatorname{cl}({\phi^{\mathrm{op}}\odot\mathrm{id}_{M_{n}}})=\operatorname{cl}({\phi^{\mathrm{op}}})\odot\mathrm{id}_{M_{n}}$ is positive. In particular, $\phi^{\mathrm{op}}$ is completely positive if and only if $\operatorname{cl}({\phi^{\mathrm{op}}})$ is completely positive. ∎
Corollary A.11.
Any completely positive unital map $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ satisfies the Kadison–Schwarz inequality: for all $x\in{}_{p}{B}$,
$\phi^{\mathrm{op}}(x^{*}x)\geq\phi^{\mathrm{op}}(x)^{*}\phi^{\mathrm{op}}(x).$
Proof.
By Proposition A.10, we can assume the map to be between C*-algebras. Now the result follows from [40, Proposition 3.3]. ∎
There are different notions of norm for the tensor product of C*-algebras. The most important ones are the minimal and the maximal C*-tensor norms. We omit a precise definition of these norms and refer the reader to the literature (see e.g. [29, Section 11.3] or [5, Section 3.3]). Despite the importance of these two norms in giving a well-defined symmetric monoidal structure (see Definition 2.31), their relevance in our current treatment is marginal.
Proposition A.12.
Let $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ be a completely positive unital map. Then $\phi^{\mathrm{op}}\odot\mathrm{id}_{{}_{p}{C}}$ is a completely positive unital map for every pre-C*-algebra ${}_{p}{C}$ with respect to both the minimal and the maximal tensor norm.
Proof.
We claim first that $\operatorname{cl}({\phi^{\mathrm{op}}\odot\mathrm{id}_{{}_{p}{C}}})\colon B\otimes C\rightsquigarrow A\otimes C$ is the extension of $\operatorname{cl}({\phi^{\mathrm{op}}})\odot\operatorname{cl}({\mathrm{id}_{{}_{p}{C}}})=\operatorname{cl}({\phi^{\mathrm{op}}})\odot\mathrm{id}_{C}$, which is completely positive by Proposition A.10. This is simple: $\operatorname{cl}({\phi^{\mathrm{op}}})\odot\mathrm{id}_{C}$ is clearly an extension of $\phi^{\mathrm{op}}\odot\mathrm{id}_{{}_{p}{C}}$ by definition, which means that the two maps must have the same extension by Lemma A.4.
We have now reduced the statement to the case of C*-algebras, where it is a standard result [5, Theorem 3.5.3]. ∎
Proposition A.13 ([40, Theorems 3.9 and 3.11]).
A positive unital map with commutative domain or commutative codomain is completely positive. In particular, all states are completely positive.
Proposition A.14 (Probabilistic Gelfand duality [22]).
The category of commutative C*-algebras and positive unital maps is dually equivalent to $\mathsf{CHausStoch}$, the category of compact Hausdorff spaces and continuous Markov kernels taking values in Radon probability measures.
Spectra of elements of a pre-C*-algebra
For a brief introduction to spectra, the reader may refer to [48, Section 1.3].
Definition A.15.
  1. (i)
    For a C*-algebra $A$ and $x\in A$, the spectrum $\mathrm{sp}\mathopen{}\mathclose{{}\left(x}\right)$ is the set of values $\lambda\in\mathbb{C}$ such that $x-\lambda 1$ is not invertible.
  2. (ii)
    For a pre-C*-algebra ${}_{p}{A}$ and $x\in{}_{p}{A}$, the spectrum $\mathrm{sp}\mathopen{}\mathclose{{}\left(x}\right)$ is the spectrum of $x$ in the completion $A$.
It is a crucial fact of the theory that every element has nonempty spectrum [28, Theorem 3.2.3].
Proposition A.16 (e.g. [9, Proposition 7.8]).
Let $x\in A$ be a self-adjoint. Then its spectrum is a compact subset of $\mathbb{R}$ with
[…]
Using Beurling’s formula [31, p. 204], we also get the following.
Corollary A.17.
The norm of a self-adjoint $x$ can be characterized in terms of states as
[…]
Corollary A.18.
Every self-adjoint $x$ satisfies $-\mathopen{}\mathclose{{}\left\lVert x}\right\rVert\leq x\leq\mathopen{}\mathclose{{}\left\lVert x}\right\rVert$. In particular, every element can be written as a linear combination of positive elements.
Corollary A.19.
Let $x\in A$. If $\phi^{\mathrm{op}}(x)=0$ for every state $\phi^{\mathrm{op}}$, then $x=0$.
Corollary A.20.
Let $\phi^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow{}_{p}{A}$ be a self-adjoint unital map which is contractive on self-adjoint elements. Then $\phi^{\mathrm{op}}$ is positive. [^21] The proof is analogous to Eric Wofsey’s answer at math.stackexchange.com/questions/3304114.
Appendix B Infinite tensor powers and permutation-invariant elements
In this appendix, we give a proof of the final sentence of Remark 3.29. Let us consider $A_{J}$, the infinite tensor power of a C*-algebra $A$ [36, Section 6.1] by some infinite set $J$. In the following, it will not matter whether we take this infinite tensor product with respect to the minimal or the maximal C*-norm. Following 3.28, every finite permutation $\sigma$ induces a bounded ***-homomorphism $A_{\sigma}^{\mathrm{op}}\colon A_{J}\rightsquigarrow A_{J}$ given by permuting the tensor factors.
Proposition B.1.
Let us consider any C*-algebra $A$, an infinite set $J$ and $x\in A_{J}$. If $A^{\mathrm{op}}_{\sigma}(x)=x$ for every finite permutation $\sigma\colon J\to J$, then $x=\lambda 1$ for some $\lambda\in\mathbb{C}$.
We thank MathOverflow users Diego Martinez and Caleb Eckhardt, whose comments were crucial for arriving at the following argument. [^22] See mathoverflow.net/questions/452405.
Proof.
By definition of $A_{J}$, there is a sequence $x_{n}\in A_{F_{n}}\subset A_{J}$, where $F_{n}$ is a finite set, such that $x=\lim_{n}x_{n}$. For clarity, let us consider $\epsilon_{n}>0$ such that $\mathopen{}\mathclose{{}\left\lVert x-x_{n}}\right\rVert\leq\epsilon_{n}$ and $\epsilon_{n}\to 0$. We pick a finite permutation $\sigma_{n}$ such that $\sigma_{n}(F_{n})\cap F_{n}=\varnothing$. Then, since $\mathopen{}\mathclose{{}\left\lVert A^{\mathrm{op}}_{\sigma_{n}}}\right\rVert\leq 1$ by Proposition A.9,
$\mathopen{}\mathclose{{}\left\lVert x_{n}-{A^{\mathrm{op}}_{\sigma_{n}}}(x_{n})}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x_{n}-x+{A^{\mathrm{op}}_{\sigma_{n}}}(x)-{A^{\mathrm{op}}_{\sigma_{n}}}(x_{n})}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x_{n}-x}\right\rVert+\mathopen{}\mathclose{{}\left\lVert A^{\mathrm{op}}_{\sigma_{n}}}\right\rVert\mathopen{}\mathclose{{}\left\lVert x-x_{n}}\right\rVert\leq 2\epsilon_{n}.$
Let us fix any state $\phi^{\mathrm{op}}\colon A_{J}\rightsquigarrow\mathbb{C}$. Our assumptions also imply
$\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}({A^{\mathrm{op}}_{\sigma_{n}}}(x_{n}))-\phi^{\mathrm{op}}(x)}\right\rVert\leq\epsilon_{n}.$
Using the completely positive map
$\psi^{\mathrm{op}}\coloneqq\phi^{\mathrm{op}}_{\mid A_{\sigma_{n}(F_{n})}}\otimes\mathrm{id}_{A_{J\setminus\sigma_{n}(F_{n})}}\colon A_{J}\rightsquigarrow A_{J\setminus\sigma_{n}(F_{n})},$
we obtain
$\mathopen{}\mathclose{{}\left\lVert x_{n}-\phi^{\mathrm{op}}({A^{\mathrm{op}}_{\sigma_{n}}}(x_{n}))1}\right\rVert=\mathopen{}\mathclose{{}\left\lVert\psi^{\mathrm{op}}(x_{n}-{A^{\mathrm{op}}_{\sigma_{n}}}(x_{n}))}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x_{n}-{A^{\mathrm{op}}_{\sigma_{n}}}(x_{n})}\right\rVert\leq 2\epsilon_{n},$
where the first equation holds because both terms can be written with tensor factors of $1$ in all tensor factors that belong to $J\setminus F_{n}$. Finally, we get
$\mathopen{}\mathclose{{}\left\lVert x-\phi^{\mathrm{op}}(x)1}\right\rVert\leq\mathopen{}\mathclose{{}\left\lVert x-x_{n}}\right\rVert+\mathopen{}\mathclose{{}\left\lVert x_{n}-\phi^{\mathrm{op}}({A^{\mathrm{op}}_{\sigma_{n}}}(x_{n}))1}\right\rVert+\mathopen{}\mathclose{{}\left\lVert\phi^{\mathrm{op}}({A^{\mathrm{op}}_{\sigma_{n}}}(x_{n}))1-\phi^{\mathrm{op}}(x)1}\right\rVert\leq 4\epsilon_{n}\to 0,$
from which we conclude that $x=\phi^{\mathrm{op}}(x)1$. ∎
Appendix C On the quantum de Finetti theorem for states
The arguments of Hulanicki and Phelps on quantum de Finetti theorems [25] are very insightful. In particular, they were able to show that the de Finetti theorem holds in a much more general context than that of C*-algebras [25, Theorem 4.1]. Their method is completely independent of Størmer’s essentially simultaneous paper [43], in which the quantum de Finetti theorem for the minimal tensor product was proven.
In this appendix, we show how the ideas of Hulanicki and Phelps that enter the proof of their Theorem 4.1 lead to a quantum de Finetti theorem for both the minimal and the maximal tensor products. Moreover, this result allows for one extra tensor factor in the style of a known criterion for the separability of states in quantum information theory [10, Theorem 1].
Lemma C.1.
In $\mathsf{CPU}_{\min/\max}$, let $\phi^{\mathrm{op}}\colon{}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}\rightsquigarrow\mathbb{C}$ be a state which is exchangeable in the first factor, and let $a\in{}_{p}{A}$ be positive. Then there exist a non-negative real number $\lambda$ and a state $\psi^{\mathrm{op}}$ exchangeable in the first factor such that the functional
$\phi^{\mathrm{op}}_{a}(y)\coloneqq\phi^{\mathrm{op}}(a\odot y),\qquad\forall\,y\in{}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}$
is simply given by $\lambda\psi^{\mathrm{op}}$.
Proof.
If $\phi^{\mathrm{op}}(a\odot 1)=0$, then for every positive $y\in{}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}$, we have
$00\leq\phi^{\mathrm{op}}(a\odot y)\leq\mathopen{}\mathclose{{}\left\lVert y}\right\rVert\phi^{\mathrm{op}}(a\odot 1)=$ 0
by Corollary A.18. This in particular implies that the statement is trivially true in this case, because $\phi^{\mathrm{op}}_{a}=0$.
If $\phi^{\mathrm{op}}(a\odot 1)\neq 0$, then this quantity is positive because $a$ is. Let us define
$\psi^{\mathrm{op}}(y)\coloneqq\frac{1}{\phi^{\mathrm{op}}(a\odot 1)}\phi^{\mathrm{op}}_{a}(y).$
By construction, $\psi^{\mathrm{op}}$ is a state and exchangeable in the first factor, therefore concluding the proof. ∎
Theorem C.2 (Quantum de Finetti, analytic version).
In $\mathsf{CPU}_{\min/\max}$, consider states ${}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}\rightsquigarrow\mathbb{C}$ that are exchangeable in the first factor. Then:
  1. (i)
    These states form a compact convex set in the weak-* topology.
  2. (ii)
    Such a state $\phi^{\mathrm{op}}\colon{}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}\rightsquigarrow\mathbb{C}$ is extremal if and only if it is of the form
    $\phi^{\mathrm{op}}=\psi^{(\mathbb{N}),\mathrm{op}}\odot\omega^{\mathrm{op}}$ (C.1)
    for $\psi^{\mathrm{op}}\colon{}_{p}{A}\rightsquigarrow\mathbb{C}$ an arbitrary state and $\omega^{\mathrm{op}}\colon{}_{p}{B}\rightsquigarrow\mathbb{C}$ a pure state.
  3. (iii)
    For every such state $\phi^{\mathrm{op}}\colon{}_{p}{A}_{\mathbb{N}}\odot{}_{p}{B}\rightsquigarrow\mathbb{C}$, there is a Radon measure $\mu$ on the product of state spaces $\Sigma A\times\Sigma B$ such that
    $\phi^{\mathrm{op}}=\int\mathopen{}\mathclose{{}\left[\psi^{(\mathbb{N}),\mathrm{op}}\odot\omega^{\mathrm{op}}}\right]\,\,\mu(\mathrm{d}\psi^{\mathrm{op}},\mathrm{d}\omega^{\mathrm{op}}).$
Proof.
Claim (i) is a direct check based upon the fact that permuting the tensor factors of ${}_{p}{A}$ is a weak-* continuous operation on the state space $\Sigma(A_{\mathbb{N}}\otimes B)$.
Concerning part (ii), note that $\phi^{\mathrm{op}}=\phi^{\mathrm{op}}_{a}+\phi^{\mathrm{op}}_{1-a}$ for any positive $a\in{}_{p}{A}$. By Lemma C.1 and extremality of $\phi^{\mathrm{op}}$, there exists a real number $\lambda\geq 0$ such that $\phi^{\mathrm{op}}_{a}=\lambda\phi^{\mathrm{op}}$. Evaluating at $1$ shows that $\lambda=\phi^{\mathrm{op}}(a\odot 1)$. Therefore,
$\phi^{\mathrm{op}}(a\odot y)=\phi^{\mathrm{op}}(a\odot 1)\,\phi^{\mathrm{op}}(y)$
for all positive $a\in{}_{p}{A}$ and $y\in{}_{p}{A_{\mathbb{N}}}\odot{}_{p}{B}$. Applying this equation inductively gives
$\begin{split}\phi^{\mathrm{op}}(a_{1}\odot\dots\odot a_{n}\odot b)&=\phi^{\mathrm{op}}(a_{1}\odot 1)\,\phi^{\mathrm{op}}(a_{2}\odot\dots\odot a_{n}\odot b)\\ &\hskip 5.20004pt\vdots\\ &=\phi^{\mathrm{op}}(a_{1}\odot 1)\cdots\phi^{\mathrm{op}}(a_{n}\odot 1)\,\phi^{\mathrm{op}}(1\odot b)\end{split}$
for all positive $a_{1},\ldots,a_{n}\in{}_{p}{A}$ and $b\in{}_{p}{B}$. We now set $\psi^{\mathrm{op}}(a)\coloneqq\phi^{\mathrm{op}}(a\odot 1)$ for all $a\in{}_{p}{A}$ and $\omega^{\mathrm{op}}(b)\coloneqq\phi^{\mathrm{op}}(1\odot b)$ for all $b\in{}_{p}{B}$. Then we already have (C.1) for simple tensors consisting of positive elements. Corollary A.18 applied to each component of a simple tensor shows that the equation holds for all simple tensors, and we then get the general case by linearity.
We now turn to part (iii). Since the map
$\Sigma A\times\Sigma B$ $\to\Sigma(A_{\mathbb{N}}\odot B)$
$(\psi^{\mathrm{op}},\omega^{\mathrm{op}})$ $\mapsto\psi^{\mathrm{op}}\odot\omega^{\mathrm{op}}$
is weak-* continuous, its image is compact and in particular closed. The claim follows now by (ii) and Choquet’s theorem. ∎
Remark C.3.
In particular, every exchangeable state on an infinite maximal tensor power of a C*-algebra $A$ can be written as an integral over product states. With an additional tensor factor $B$, this also generalizes the separability criterion of [10, Theorem 1] to the infinite-dimensional case with maximal tensor product.